### v1

# Transformer Alignment in Large Language Models

Abstract
Large Language Models (LLMs) have made significant strides in natural language processing, and a precise understanding of the internal mechanisms driving their success is essential. We regard LLMs as transforming embeddings via a discrete, coupled, nonlinear, dynamical system in high dimensions. This perspective motivates tracing the trajectories of individual tokens as they pass through transformer blocks, and linearizing the system along these trajectories through their Jacobian matrices. In our analysis of 38 openly available LLMs, we uncover the alignment of top left and right singular vectors of Residual Jacobians, as well as the emergence of linearity and layer-wise exponential growth. Notably, we discover that increased alignment positively correlates with model performance. Metrics evaluated post-training show significant improvement in comparison to measurements made with randomly initialized weights, highlighting the significant effects of training in transformers. These findings reveal a remarkable level of regularity that has previously been overlooked, reinforcing the dynamical interpretation and paving the way for deeper understanding and optimization of LLM architectures.
1 Introduction
Large language models (LLMs), as exemplified by BERT and GPT-series [Devlin et al., 2019, Brown et al., 2020], have revolutionized the field of natural language processing through their adoption of the transformer architecture [Vaswani et al., 2017]. Despite their widespread success, the internal mechanisms that underpin their performance are not fully understood.
Previous works viewed certain types of deep networks as implementing discrete, nonlinear dynamical systems, operating in high dimensions [Greff et al., 2016, Papyan et al., 2017, Ebski et al., 2018, Chen et al., 2018, Bai et al., 2019, Rothauge et al., 2019, Li and Papyan, 2023, Gai and Zhang, 2021, Haber and Ruthotto, 2017, Ee, 2017]. The term discrete reflects the network’s finite depth; nonlinear refers to the model’s nonlinear components; and dynamical is due to the residual connections spanning various layers. This motivates us to trace the dynamics of individual tokens as they traverse through the numerous transformer blocks, and to linearize the system through its Jacobian matrices.
1.1 Residual Alignment
Our investigation draws inspiration from a recent study by Li and Papyan [2023] on Residual Networks (ResNets) [He et al., 2016], which uncovered a phenomenon they termed Residual Alignment (RA), marked by several distinct characteristics in hidden representations of ResNets:
  • (RA1)
    Linear and equispaced trajectories in layer-wise progression of intermediate representations.
  • (RA2)
    Alignment of top left and right singular vectors in the linearizations of residual blocks.
  • (RA3)
    The rank of Residual Jacobians is at most the number of classes.
  • (RA4)
    Top singular values of Residual Jacobians scale inversely with depth.
1.2 Contributions
We uncover Transformer Alignment (TA), characterized by two properties (analogous to those of RA) that consistently emerge with training in 38 openly available base and fine-tuned LLMs (Appendix A.1) including the Falcon [Almazrouei et al., 2023], Llama-3 [AI@Meta, 2024], Llama-2 [Touvron et al., 2023], MPT [Team, 2023a, b], Mistral v0.1 [Jiang et al., 2023], Gemma [Team et al., 2024], NeoX [Black et al., 2022], Neo [Black et al., 2021, Gao et al., 2020], and Pythia [Biderman et al., 2023] families. The properties are given by
  • (TA1)
    Emergence of linearity and exponentially distanced trajectories of intermediate representations in layer-wise progression.
  • (TA2)
    Alignment of top left and right singular vectors in the linearizations of transformer blocks through their Residual Jacobians.
Due to the large vocabulary size relative to the embedding size in LLMs, Residual Jacobians maintain rank of at most the vocabulary size, and so the previous (RA3) immediately holds. Further, trained LLMs do not exhibit a consistent inverse scaling of top singular values with depth, and scaling varies considerably between models, demonstrating a lack of (RA4). In contrast to ResNets, the layer-wise progression of transformer representations is better described with an exponential evolution, rather than an equispaced linear progression.
Additionally, we explore the relationship between (TA2) and generalization, and uncover that increased alignment of Residual Jacobian singular vectors correlates with improved Open LLM leaderboard [Beeching et al., 2023] benchmark score (Figure 9), raising valuable insight to the mathematical properties of LLM transformers and their connections to generalization.
2 Background on Large Language Models
In the input layer, $l=0$, textual prompts undergo tokenization and are combined with positional encodings to create an initial high-dimensional embedding, denoted by $x_{i}^{0}\in\mathbb{R}^{d_{\text{model}}}$ for the $i^{\text{th}}$ token. When these embeddings are stacked together, they form a matrix:
$X^{0}=(x_{1}^{0},x_{2}^{0},\ldots,x_{n}^{0})\in\mathbb{R}^{d_{\text{model}}\times n}.$ (1)
The embeddings then pass through $L$ transformer blocks:
$X^{0}\xrightarrow{f_{\text{block}}^{1}}X^{1}\xrightarrow{f_{\text{block}}^{2}}\cdots X^{L-1}\xrightarrow{f_{\text{block}}^{L}}X^{L}.$ (2)
Here, $X^{l}=f_{\text{block}}^{l}(X^{l-1})$ denotes the embeddings after the $l^{\text{th}}$ block, consisting of causal multi-headed attention (MHA), a feed-forward network (FFN), and normalization layers (LN) with residual connections:
$h^{l+1}(X^{l})$ $=\text{MHA}(\text{LN}(X^{l}))$ (3)
$g^{l+1}(X^{l})$ $=\text{LN}(X^{l}+h^{l+1}(X^{l}))$ (4)
$f_{\text{block}}^{l+1}(X^{l})$ $=X^{l}+h^{l+1}(X^{l})+\text{FFN}(g^{l+1}(X^{l})),$ (5)
where the MHA, LN, FFN are implicitly indexed by layer. In the final representation, an additional layer normalization is applied:
$f_{\text{block}}^{L}(X^{L-1})$ $=\text{LN}(X^{L-1}+h^{L}(X^{L-1})+\text{FFN}(g^{L}(X^{L-1}))).$ (6)
The output $X^{L}$ from the final block $f^{L}$ is passed into a bias-free linear layer $M\in\mathbb{R}^{d_{\text{vocab}}\times d_{\text{model}}}$, with $d_{\text{vocab}}$ denoting the size of the token vocabulary and $d_{\text{model}}$ is the dimension of the token embeddings. This layer $M$ computes final-layer logits for each token embedding, $\ell_{i}=Mx_{i}^{L}$. The prediction for the next token is then determined by selecting the maximal logit value: $\arg\max_{v\in\text{tokens}}\ell_{v,n}.$
3 Methods
3.1 Suite of Large Language Models
Our study evaluates a total of 38 LLMs (24 base LLMs and 14 fine-tuned, see Appendix A.1) that were independently trained by various individuals and organizations. These models, provided through HuggingFace [Wolf et al., 2020], vary in terms of parameter budgets, number of layers, hidden dimensions, and training tokens. A summary of the models under consideration is presented in Table 1 of Appendix A.1. Measurements are made for models with varying parameter sizes with GPUs satisfying appropriate memory requirements, further detailed in Appendix A.3.
3.2 Prompt Data
We evaluate these LLMs using prompts of varying length, ambiguity, and context, sourced from the test set of ARC [Clark et al., 2018], GSM8K [Cobbe et al., 2021], HellaSwag [Zellers et al., 2019], MMLU [Hendrycks et al., 2021], Truthful QA [Lin et al., 2022], and Winogrande [Sakaguchi et al., 2019]. These datasets set the performance benchmarks on the HuggingFace Open LLM Leaderboard [Beeching et al., 2023] since they encompass a diverse set of language tasks. Post-prompting, we assess the models using several metrics, detailed in the following subsections.
3.3 Alignment of Singular Vectors of Residual Jacobians
To demonstrate the presence of (TA2), we examine the properties of the transformer blocks and the relationships between them. This is done by analyzing the linearizations of the blocks given by their Residual Jacobian matrices
$J_{l}^{i}=\frac{\partial}{\partial x_{i}^{l-1}}\left(h^{l}(X^{l-1})+\text{FFN\textsubscript{$l$}}(g^{l}(X^{l-1}))\right)_{i},$ (7)
defined for each $l=1,\ldots,L$, and $i=1,\ldots,n$. Note that this is the Jacobian matrix for each transformer block without the contribution from the skip connection from the input of the block, analogous to the quantities measured by Li and Papyan [2023].
The singular value decompositions of these $J_{l}^{i}$ are computed, i.e., $J_{l}=U_{l}S_{l}V_{l}^{\top}$ (with superscript $i$, indicating the token, omitted for clarity), where $U_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ and $V_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ are the matrices of left and right singular vectors respectively, and $S_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ is the singular value matrix.
The matrices $A_{i,j,K}\coloneqq U_{j,K}^{\top}J_{i}V_{j,K}$ are plotted over all pairs of depths $i,j\in\{1,\ldots,L\}$, where $U_{j,K}$ and $V_{j,K}$ are the sub-matrices with columns that are the top-$K$ left and right singular vectors of $J_{j}$, respectively.
To quantify the misalignment of singular vectors of the Residual Jacobians at depths $i$ and $j$, we measure how far the $A_{i,j,K}$ are from being diagonal. More precisely, we define the misalignment of $A_{i,j,K}$ as
$m(A_{i,j,K})=\lVert A_{i,j,K}-\text{Diag}(A_{i,j,K})\rVert_{F},$ (8)
where $\text{Diag}(A_{i,j,K})$ is the matrix $A_{i,j,K}$ with all non-diagonal entries replaced by zero and $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm.
3.4 Layer-Wise Exponential Growth
To demonstrate the presence of (TA1), we measure the expodistance of the hidden trajectories, defined below. In this context, each layer in the model corresponds to a unit time step of a dynamical system, so we denote $\dot{x}_{i}(t)=x_{i}(t+1)-x_{i}(t)$ the time derivative of the $i^{\text{th}}$ token. Assuming exponential growth of the embedding norms as they flow through the hidden layers, we can write,
$\|\dot{x}_{i}(t)\|$ $\approx e^{\alpha t}\|\dot{x}_{i}(0)\|$ (9)
for some fixed $\alpha\in\mathbb{R}$. We quantify the validity of the approximation in Equation (9) by measuring the coefficient of variation of $\alpha$ for each trajectory. Namely, following (9),
$\alpha_{t}^{i}\approx\frac{1}{t}\ln\left(\frac{\|\dot{x}_{i}(t)\|}{\|\dot{x}_{i}(0)\|}\right)$ (10)
for each time (layer) $t$ and token $i$. Under exponential growth, it is expected that $\alpha_{t}^{i}$ is independent of time. We then denote the expodistance (ED) of the trajectory of the $i^{th}$ token of a given sequence by
$ED_{i}=\frac{\text{Var}_{t}\alpha_{t}^{i}}{(\text{Avg}_{t}\alpha_{t}^{i})^{2}}$ (11)
This measurement is motivated by the parametrization discussed in Appendix A.2 and serves as a method to test the validity of the linearization presented in Equation (15).
3.5 Linearity of Trajectories
Linearity in intermediate embeddings (TA1) is quantified with the line-shape score (LSS), defined by Gai and Zhang [2021] as
$\text{LSS}_{i}^{0,\ldots,L}=\frac{L}{\left|\left|\tilde{x}^{L}_{i}-\tilde{x}^{0}_{i}\right|\right|_{2}},$ (12)
where $\tilde{x}_{i}^{0}=x_{i}^{0}$, i.e., the input embeddings passed to the LLM, and $\tilde{x}_{i}^{l}$ is defined recursively as
$\tilde{x}^{l}_{i}=\tilde{x}^{l-1}_{i}+\frac{x^{l}_{i}-x^{l-1}_{i}}{\left|\left|x^{l}_{i}-x^{l-1}_{i}\right|\right|_{2}}\ \ \mbox{for}\ \ l=1,\ldots,L.$ (13)
Note that $\text{LSS}\geq 1$, with $\text{LSS}=1$ if and only if the intermediate representations $x_{i}^{0},\ldots,x_{i}^{L}$ form a co-linear trajectory.
3.6 Visualization of Trajectories with PCA
Each token, with initial embedding $x_{i}^{0}$, forms a trajectory $x_{i}^{0},x_{i}^{1},\ldots,x_{i}^{L}$ as it passes through the $L$ transformer blocks. The dynamics in high-dimensional space are visualized through a 2-dimensional principal component (PC) projection, $\text{PC}_{L}$, fitted to the last layer embeddings $X^{L}=(x_{1}^{L},x_{2}^{L},\ldots x_{n}^{L})$. The projected embeddings, $\text{PC}_{L}(x_{i}^{0}),\,\text{PC}_{L}(x_{i}^{1}),\,\ldots,\,\text{PC}_{L}(x_{i}^{L})$, are plotted for each of the $i=1,\ldots,n$ trajectories.
4 Results
4.1 Alignment of Singular Vectors of Residual Jacobians
In trained LLMs, we observe (TA2), that is, alignment of the top singular vectors of the Jacobians $J_{l}$ across depth (Figure LABEL:fig:paris_all_8to16 bottom row), evident in the distinct diagonal lines present in the matrix subplots. This phenomenon is consistently observed across LLMs, similar to (RA2), which was observed by Li and Papyan [2023] in ResNets. On the other hand, in untrained models (Figure LABEL:fig:paris_all_8to16 top row), there is no alignment between singular vectors of Residual Jacobians across different depths. There is alignment along the diagonal, however, because each Residual Jacobian is trivially diagonalized by its own singular vectors in $A_{i,j,K}$. This suggests that (TA2), the alignment of singular vectors of Residual Jacobians, emerges through training.
To understand this phenomenon more precisely, for each LLM, we measure the average alignment (which we define to be negative misalignment) across prompts in the 6 evaluation datasets (Section 3.2), where for each prompt, the misalignment $m(A_{i,j,K})$ is averaged over layers $i,j=1,\ldots,L$. We plot the negative misalignment values against the benchmark scores across several LLMs (Figure 9). Our results reveal a consistent pattern: an increase in alignment corresponds to an increase in benchmark scores. This observation suggests a compelling relationship between stronger alignment of singular vectors of Residual Jacobians and improved generalization performance.
Figure 9: Relationship between Alignment and Benchmark Score. The plot illustrates the relationship between average alignment (negative misalignment) and benchmark score across various LLMs. The left subplot includes models fine-tuned for coding and instruction, while the right excludes these models. We make this distinction as these fine-tuned models are for code and instruction, rather than being fine-tuned on the benchmark tasks. The data shoes a noticeable trend: increased alignment correlates with elevated benchmark scores. In particular, when excluding the code and instruction models, the linear regression gives an $R^{2}$ value of $0.7214$. This suggests that stronger singular vector alignment may correspond with enhanced generalization performance.
4.2 Linearity of Trajectories
We observe (TA1) in each LLM, evident in the LSS of layer-wise trajectories at initialization and after training evaluated over benchmark prompts (Figure 10), in agreement with (RA1) in ResNets. Linear and expansive dynamical behavior is demonstrated in Llama-3 70B, Mistral 30B, and NeoX 20B through low dimensional projections of embedding trajectories (Figure LABEL:figs:pca). The LSS of the trajectories has an average value $4.25$ across trained models, while taking average values of $6.54$ at initialization, and is evidenced by the low variation across benchmark prompts (Figure 10).
Figure 10: (TA1): Linearity of layer-wise trajectories. The figure depicts the line-shape score (LSS) of embedding trajectories, as discussed in Section 4.2), computed on 1,200 prompts of the HuggingFace Open LLM Leaderboard (Section 3.2) for a variety of trained (black) and randomly initialized (red) LLMs (Appendix A.1). Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model. Models are sorted by increasing benchmark performance. Across every model, training increases linearity (lower LSS) of embedding trajectories when compared to that of untrained models (greater LSS).
After training, LSS appears largely independent of the parameter budget, embedding dimension, and the HuggingFace OpenLLM benchmark score. At initialization, the non-linearity of the trajectories grows with model size, embedding dimension, and number of layers (Figure 10). Moreover, the discrepancy between trained and untrained models increases with the number of parameters, hidden layers, and embedding dimension.
4.3 Layer-wise Growth
Across all LLMs considered, the hidden trajectories exhibit distinctly exponential growth that emerges with training (Figure 11), displaying the second property of (TA1). This property, which arises as a result of Jacobian alignment, is distinct from (RA1) in the sense that distances between consecutive layers grow exponentially instead of linearly with depth, and highlights a key difference between the trajectories in ResNet classifiers and transformers. Most of the LLMs considered exhibit a low coefficient of variation across prompts, showing the robustness of this property across a variety of tasks. The expodistance measurements show little dependence on model parameters, and take on consistent and small values across the models (Figure 11).
Figure 11: (TA1): Coefficient of variation of layer-wise exponent. The figure presents the coefficient of variation of the layer-wise exponent (Section 3.4) computed on 1,200 prompts of the HuggingFace Open LLM Leaderboard (Section 3.2) for trained LLMs (Appendix A.1). Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model. LLMs are sorted by increasing benchmark performance, while the vertical axis has a base 10 logarithmic scale. The coefficient of variation for most models is low, suggesting that the growth of the embedding norms follows an exponential pattern across different layers, characterized by a relatively constant exponent.
At initialization, each model consistently displays equally (rather than exponentially) distanced trajectories, as reflected by the extremely low coefficients of variation in the norms of their layer-wise differences (Figure 23). Conversely, the trajectories of the trained models appear to be significantly less equidistant across layers, exhibiting substantially larger coefficients of variation with far less consistency across prompts. This behaviour, however, is expected and is discussed in more detail in Section 5.4.
5 Discussion
5.1 Correlation with Generalization
As discussed in Section 4.1, stronger alignment of Residual Jacobians singular vectors is related to improved generalization (Figure 9). This relationship between (TA2) with LLM benchmark scores may assist in evaluating the training of transformers. Understanding the relationship between training methods and regularity in representations holds potential implications for the enhancement of LLMs. This finding suggests that development of techniques to amplify Transformer Alignment may lead to favourable model performance. The connection between hidden representation regularity, generalization, and LLM training, as prompted by our work, presents an intriguing research direction with important practical implications.
5.2 Regularity and Training
Many LLMs are based on the same transformer architecture, yet when trained by independent individuals and organizations, training directly affects the emergence of residual Jacobian alignment (Figure 9), presence of linearity (Figure 10), and extent of exponential growth (Figure 11). The effect of training is further evident when comparing base LLMs to their fine-tuned counterparts, since differences in measurements are a direct product of further optimization of the same weights. Our findings further highlight the important effect of training on the dynamical properties of hidden trajectories in transformers.
5.3 Comparison to ResNets
Linearity and Equidistance.
Linearity in hidden trajectories, as observed in ResNets (RA1) Li and Papyan [2023], Gai and Zhang [2021], also emerges with training in LLMs (TA1). The mean LSS value among the evaluated LLMs is $4.24$ (Figure 10), greater than LSS measurements observed for ResNets (Gai and Zhang [2021], page 18) which range between $2.0$-$3.0$. In both architectures, training induces improved linearity and regularity (Figure LABEL:figs:pca) in trajectories. In contrast to ResNets, trajectories are not equidistant, instead showing exponential growth between layers (Figure 11). A low coefficient of variation displays the presence of (TA1) on token trajectories, in analogy with the equidistant spacing of (RA1). In both ResNet classifiers and LLM transformers, there is an evident level of regularity in hidden representations (Figure LABEL:figs:pca).
Residual Jacobian Singular Vector Alignment.
The alignment of singular vectors of Residual Jacobians is present across LLMs (Figure 9), similar to how it occurs in ResNets [Li and Papyan, 2023], and does not occur at initialization.
Rank of Residual Jacobians.
Li and Papyan [2023] show that residual Jacobians have rank at most $C$, the number of classes. This analogous result automatically holds for LLMs since the vocabulary size is significantly greater than the embedding dimension of the transformer blocks.
Singular Value Scaling.
Li and Papyan [2023] observe that top singular values of residual Jacobians scale inversely with depth. In trained LLMs, however, top singular values do not show a consistent depth scaling across models (Figure LABEL:fig:svalues_trained), notably differing from classification ResNets. In addition, the distribution of singular values at each layer varies significantly between models. Interestingly, singular value scaling is actually more present in untrained transformers (FigureLABEL:fig:svalues_untrained).
5.4 Emergence of Exponential Growth
Trained ResNets distinctly feature equidistance in hidden trajectories, while measurements in transformers display a layer-wise exponential growth in norm. [Li and Papyan, 2023] prove that equidistance in certain classifiers emerges under the assumptions of (RA 2-4), in particular, the authors assume that top singular values of residual Jacobians scale inversely with model depth. In transformers, top singular values do not show a similar scaling, possibly suggesting that (RA4) is a necessary condition for equidistance.
Emergence of (TA1) appears to be a result of (TA2) and a relationship between the singular values, lying analogously to the emergence of (RA1) as a result of (RA 2-4). (TA2) implies that linearizations of hidden trajectories take a simple form after training, satisfying a linear difference equation (see Appendix A.2). In contrast to (RA4), the top singular values of the Jacobian blocks exhibit low variation across depth in some models (Figure LABEL:fig:svalues_trained). Under these assumptions, decomposing the matrix $A$ in Equation (16) as $A=U\Sigma V^{T}$, we obtain
$|x(t)|\approx\exp(\sigma t)|x(0)|,$ (14)
where $\sigma$ is the top singular value. Since $\sigma$ is independent of depth, the above norm will be exponential in $t$, and directly explains the pattern seen in Figure 11.
Common in many models at initialization, the top singular values of the Jacobian matrices decay reciprocally with depth (Figure LABEL:fig:svalues_untrained). As a result, relatively linear trajectories appearing in the untrained models is not surprising. Randomly initialized Jacobian matrices with decaying singular values, coupled with Equation 17 results in the approximate exponential growth becoming linearized due to the receding impact of each layer.
6 Related Work
Residual Networks. ResNets have been viewed as an ensemble of shallow networks [Veit et al., 2016], with studies delving into the scaling behaviour of their trained weights [Cohen et al., 2021]. The linearization of residual blocks by their Residual Jacobians was first explored by Rothauge et al. [2019], who examined Residual Jacobians and their spectra in the context of stability analysis, and later by Li and Papyan [2023] who discovered Residual Alignment. We continue this line of work by investigating Residual Jacobians in transformer architectures and uncover a similar phenomenon in pretrained LLMs.
In-context learning. Previous works have observed that LLMs are able to perform in context learning [von Oswald et al., 2023, Bai et al., 2024, Ahn et al., 2023, Akyürek et al., 2023, Xie et al., 2021, Hahn and Goyal, 2023, Xing et al., 2024]: the model is capable of performing a particular machine learning task through a few examples provided in a single prompt or context. Subsequent theoretical works explained this phenomenon by postulating that intermediate representations of deep networks are capable of emulating classical learning algorithms like solving least squares through gradient descent, solving LASSO, and even training a shallow network.
Our work provides empirical evidence for these works by demonstrating that the transformer blocks of LLMs are coupled in a way that would only emerge if consecutive blocks jointly try to solve a specific problem.
Neural ODE. Introduced by [Chen et al., 2018], Neural ODE view ResNets as a discretized dynamical process, with past work [Sander et al., 2022] showing the convergence of Residual Networks to a system of linear ODE. [Zhong et al., 2022, Li et al., 2021] studied neural ODE specific perspective to transformers and other LLM architectures. In our work, the presence of residual connections and emergence of Jacobian Alignment in transformers suggests that a discretization of a simple ODE emerges in LLMs.
Information Processing in LLMs. Previously, Katz and Belinkov [2023], Bietti et al. [2023] have described memory and semantic flow in intermediate states of transformers while van Aken et al. [2019] have examined BERT hidden trajectories. Investigation of LLMs for reasoning tasks and interpretability remains an important research focus [Huang and Chang, 2023].
Hidden Representation Dynamics. Previous studies [Gai and Zhang, 2021, Haber and Ruthotto, 2017, Ee, 2017], have focused on interpreting deep neural networks from a dynamical systems perspective. In particular, the layer-wise transformations between transformer blocks are interpreted as discrete approximations to continuous curves through representation space. Throughout training, numerical experiments have shown that both plain networks and ResNets are aptly approximated by a geodesic curve in Wasserstein space [Gai and Zhang, 2021]. Other interpretations involve partitioning activation space into basins of attraction, with fixed points representing discrete thoughts [Nam et al., 2023]. For further related works, see [Geshkovski et al., 2023b, a, Tarzanagh et al., 2023, Valeriani et al., 2023].
Structure in Hidden Representations. Emergence of regularity in last-layer representations during training was noticed in Neural Collapse [Papyan et al., 2020]. Motivated by this phenomenon, subsequent works examined structure of hidden representations and presented theoretical results describing the emergence of Neural Collapse [Wang et al., 2024a, Parker et al., 2023, Zangrando et al., 2024, Garrod and Keating, 2024, Wang et al., 2024b, Hoyt and Owen, 2021, Arous et al., 2023, Zarka et al., 2021, Ben-Shaul and Dekel, 2022, Papyan, 2020, Súkeník et al., 2023]. Our work studies the representations of all layers in LLMs in a local sense through individual trajectories, and uncovers regularity in the representations in LLMs.
7 Limitations
A main limitation of the scope of our analysis is that due to the significant compute resources required to train LLMs of the size studied in this paper, our exploration is restricted to pretrained LLMs and those fine-tuned on them. Several of these models are trained independently of each other, without carefully controlling for different experimental parameters. Consequently, direct comparisons between models become challenging, and so the cause of varying levels of regularity between LLMs remains unclear. Nonetheless, this makes the consistent emergence of these properties interesting for further study.
8 Conclusion
A detailed description of the complexities of transformer blocks in the context of language modelling is an active area of research. Our primary goal was to enhance the understanding of the mechanics underlying transformer architectures. We provided a quantitative description of embedding trajectories within LLMs. In a large collection of models, independently trained by various organizations, we discover a remarkable level of regularity among transformer representations characterized by the alignment of residual Jacobians and emergence of linear and exponentially growing trajectories. We observe a correlation between the strength of alignment of residual Jacobian singular vectors and LLM benchmark scores, highlighting the significance of (TA2) and it’s implications regarding performance. These findings open avenues for future research for a deeper understanding the connections between regularity of hidden representations and model specifications, LLM architecture, and generalization.
Appendix A Appendix and Supplementary material
A.1 Suite of Large Language Models and Prompt Data
We evaluate 38 total LLMs (24 base LLMs and 14 fine-tuned) on 6 datasets from the HuggingFace Open LLM leaderboard [Beeching et al., 2023].
Base models.
Falcon (40B, 7B) [Almazrouei et al., 2023], Llama-3 (70B, 8B) [AI@Meta, 2024], Llama-2 (70B, 13B, 7B) [Touvron et al., 2023], MPT (30B, 7B) [Team, 2023a, b], Mistral v0.1 [Jiang et al., 2023], Gemma (7B, 2B) [Team et al., 2024], Gemma 1.1 (7B, 2B), NeoX (20B) [Black et al., 2022], Neo (2.7B) [Black et al., 2021, Gao et al., 2020], Pythia (6.9B).[Biderman et al., 2023], and GPT-2 (1.5B, 774M, 355M, 117M) [Radford et al., 2019].
Fine-tuned models.
CodeLlama (34B, 13B, 7B) [Rozière et al., 2024], CodeLlama Instruct (34B, 13B, 7B) [Rozière et al., 2024], Mistral-v0.1 Instruct (7.3B) [Jiang et al., 2023], Mistral-v0.2 Instruct [Jiang et al., 2023], CodeGemma [Team et al., 2024].
Datasets.
ARC [Clark et al., 2018], HellaSwag [Zellers et al., 2019], MMLU [Hendrycks et al., 2021], Truthful QA [Lin et al., 2022], WinoGrande [Sakaguchi et al., 2019].
Table 1: LLMs featured in the experiments throughout paper. Included in the table is the parameter budget of each model, the embedding dimension, the number of training tokens, and the Open LLM leaderboard [Beeching et al., 2023] benchmark score.
Model Parameters Layers ($L$) Dim. ($d_{\text{model}}$) Training Tokens Score
Llama-3 70 B 80 8192 15 T
8 B 32 4096 15 T 62.35
Llama-2 70 B 80 8192 2 T 66.05
13 B 40 5120 2 T 55.69
7 B 32 4096 2 T 50.97
Codellama 34 B 48 8192 2 T 55.33
13 B 40 5120 2 T 45.82
7 B 32 4096 2 T 39.81
Codellama (IT) 34 B 48 8192 2 T 43.0
13 B 40 5120 2 T 37.52
7 B 32 4096 2 T 40.05
Orca 13 B 40 5120 2 T 58.64
7 B 32 4096 2 T 54.55
Falcon 40 B 60 8192 1 T 58.07
7 B 32 4544 1.5 T 44.17
Falcon (IT) 40 B 60 8192 1 T 43.26
7 B 32 4544 1.5 T
MPT 30 B 48 7168 1 T 66.98
7 B 32 4096 1 T 56.83
Phi 2 B 32 2560 1.4 T 61.33
1.5 B 24 2048 150 B 47.69
1 B 24 2048 54 B
Mistral-v0.1 7.3 B 32 4096 60.97
Mistral-v0.1 (IT) 7.3 B 32 4096 54.96
Mistral-v0.2 (IT) 7.3 B 32 4096 65.71
Gemma 7 B 28 3072 6 T 64.29
2 B 18 2048 6 T 42.75
Gemma-1.1 7 B 28 3072 6 T 30.0
2 B 18 2048 6 T 60.09
Codegemma 7 B 28 3072 6 T 56.73
2 B 18 2048 6 T 32.19
Neo 20 B 44 6144 41.69
2.7 M 32 2560 0.42 36.20
Pythia 6.9 M 32 4096 0.3 T 39.30
GPT-2 1.5 B 48 1600 34.12
774 M 36 1280 32.07
355 M 24 1024 29.87
117 M 12 768 28.53
A.2 Dynamical Motivation
The equality of top left and right singular vectors (RA2) suggests that the linearizations form a simple dynamical system that acts on representations. Consider a homogenous linear ODE:
$x^{\prime}=Ax$ (15)
Its solution is given by
$x(t)=\exp(tA)x(0)=\lim_{L\to\infty}\left(I+\frac{At}{L}\right)^{L}x(0)$ (16)
where the product is repeated $L$ times. Expanding the brackets shows that $x(t)$ can be thought of as a collection of many paths of various lengths, due to the binomial identity. This agrees with Veit et al. [2016] which views ResNets as
$x(t)=\left(I+\frac{tA_{1}}{L}\right)\left(I+\frac{tA_{2}}{L}\right)\dots\left(I+\frac{tA_{L}}{L}\right)x(0)$ (17)
However, Veit et al. [2016] do not make any assumptions about the alignment of the various $A_{i}$ matrices. The Residual Alignment phenomenon suggests ResNets as implementing the simpler system
$x(t)=\left(I+\frac{tA}{L}\right)\left(I+\frac{tA}{L}\right)\dots\left(I+\frac{tA}{L}\right)x(0),$ (18)
where all the $A_{i}$ matrices are aligned. One benefit of this interpretation is that, as $L\to\infty$, we can write $x(t)$ in a simple closed form, namely $x(t)=\exp(tA)x(0)$.
Borrowing from the dynamical perspective, we quantify the similarity of hidden trajectories to the evolution of a linear ODE, in order to detect the emergence of a simple linearization to representations.
A.3 Experimental and Implementation Details
The source code used to produce the results reported in this experiment has been included as supplemental material. Models with varying parameter sizes are loaded on GPUs with appropriate memory requirements: NVIDIA A40 ($n_{\text{param}}\geq 40B$), NVIDIA Quadro RTX 6000 for Gemma variants and when ($40B>n_{\text{param}}>13B$), and NVIDIA Tesla T4 when ($13B\geq n_{\text{param}}$) except Gemma variants. 1,200 prompts from the OpenLLM leaderboard were evaluated in variable batch sizes were queued on a SLURM cluster, with appropriate adjustments depending on the memory required to load the LLM.
  • $13B\geq n_{\text{param}}$: 100 prompts per batch, except Gemma variants, which used 25 prompts per batch. The larger memory requirement for Gemma variants is likely due to the much larger vocabulary size in the model.
  • $40B>n_{\text{param}}>13B$: 10 prompts per batch, except NeoX 20B which used 100 prompts per batch.
  • $n_{\text{param}}\geq 40B$: 50 prompts per batch.
Due to the high memory requirement for computing Residual Jacobians, for experiments involving Residual Jacobians, NVIDIA Quadro RTX 6000 was used additionally for $13B>n_{\text{param}}\geq 7B$ and corresponding models were quantized. Additionally, due to compute restrictions, alignment of singular vectors of Residual Jacobians was computed on a smaller subset of the 1,200 prompts.
A.4 Additional Experimental Results
Figure 23: Coefficient of variation of layer-wise equidistance. Variation of layer-wise equidistance (Section 3.4) computed over 1,200 prompts from the HuggingFace Open LLM Leaderboard datasets (Section 3.2) on a suite of untrained LLMs (Appendix A.1). Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model. The models are sorted by increasing benchmark performance.

### v2

@@
-Transformer Alignment in Large Language Models
+Transformer Block Coupling and its Correlation with Generalization in LLMs
 Abstract
 Large Language Models (LLMs) have made significant strides in natural language processing, and a precise understanding of the internal mechanisms driving their success is essential.
-We regard LLMs as transforming embeddings via a discrete, coupled, nonlinear, dynamical system in high dimensions.
-This perspective motivates tracing the trajectories of individual tokens as they pass through transformer blocks, and linearizing the system along these trajectories through their Jacobian matrices.
-In our analysis of 38 openly available LLMs, we uncover the alignment of top left and right singular vectors of Residual Jacobians, as well as the emergence of linearity and layer-wise exponential growth.
-Notably, we discover that increased alignment positively correlates with model performance.
-Metrics evaluated post-training show significant improvement in comparison to measurements made with randomly initialized weights, highlighting the significant effects of training in transformers.
-These findings reveal a remarkable level of regularity that has previously been overlooked, reinforcing the dynamical interpretation and paving the way for deeper understanding and optimization of LLM architectures.
+In this work, we trace the trajectories of individual tokens as they pass through transformer blocks, and linearize the system along these trajectories through their Jacobian matrices.
+By examining the relationships between these Jacobians, we uncover a transformer block coupling phenomenon in a variety of LLMs, characterized by the coupling of their top singular vectors across tokens and depth.
+Our findings reveal that coupling positively correlates with model performance, and that this relationship is stronger than with other hyperparameters, namely parameter budget, model depth, and embedding dimension.
+We further investigate the emergence of these properties through training, noting the development of coupling, as well as an increase in linearity and layer-wise exponential growth in the token trajectories.
+These collective insights provide a novel perspective on the interactions between token embeddings, and prompt further approaches to study training and generalization in LLMs.
 1 Introduction
-Large language models (LLMs), as exemplified by BERT and GPT-series [Devlin et al., 2019, Brown et al., 2020], have revolutionized the field of natural language processing through their adoption of the transformer architecture [Vaswani et al., 2017].
-Despite their widespread success, the internal mechanisms that underpin their performance are not fully understood.
-Previous works viewed certain types of deep networks as implementing discrete, nonlinear dynamical systems, operating in high dimensions [Greff et al., 2016, Papyan et al., 2017, Ebski et al., 2018, Chen et al., 2018, Bai et al., 2019, Rothauge et al., 2019, Li and Papyan, 2023, Gai and Zhang, 2021, Haber and Ruthotto, 2017, Ee, 2017].
-The term discrete reflects the network’s finite depth; nonlinear refers to the model’s nonlinear components; and dynamical is due to the residual connections spanning various layers.
-This motivates us to trace the dynamics of individual tokens as they traverse through the numerous transformer blocks, and to linearize the system through its Jacobian matrices.
-1.1 Residual Alignment
-Our investigation draws inspiration from a recent study by Li and Papyan [2023] on Residual Networks (ResNets) [He et al., 2016], which uncovered a phenomenon they termed Residual Alignment (RA), marked by several distinct characteristics in hidden representations of ResNets:
-  • (RA1)
-    Linear and equispaced trajectories in layer-wise progression of intermediate representations.
-  • (RA2)
-    Alignment of top left and right singular vectors in the linearizations of residual blocks.
-  • (RA3)
-    The rank of Residual Jacobians is at most the number of classes.
-  • (RA4)
-    Top singular values of Residual Jacobians scale inversely with depth.
+In recent years, many openly available Large Language Models (LLMs) have been released, achieving state-of-the-art results on task-specific benchmarks.
+The abundance of models, each with differing architecture and training methodology, motivates comparing the underlying mechanisms that drive generalization.
+Transformers (Vaswani et al., 2017) can be represented as discrete, nonlinear, coupled dynamical systems, operating in high dimensions (Greff et al., 2016; Papyan et al., 2017; Haber & Ruthotto, 2017; Ee, 2017; Ebski et al., 2018; Chen et al., 2018; Bai et al., 2019; Rothauge et al., 2019; Gai & Zhang, 2021; Li & Papyan, 2023).
+The term discrete reflects the network’s finite depth; nonlinear refers to the model’s nonlinear components; coupled refers to the interdependent token trajectories enabled by self-attention, and dynamical is due to the residual connections spanning various layers.
+Throughout this work, we investigate whether we can identify structural characteristics that consistently occur in pretrained LLMs that are related to their generalization, as measured on the HuggingFace OpenLLM Leaderboard benchmark.
+To this end, we study and quantify the trajectories of token embeddings and their interactions, through linearizations of LLM layers given by the Jacobians of the associated transformer blocks.
+This perspective raises several questions:
+Q1.
+What regularity properties do these trajectories exhibit individually, and what are their relations with one another?
+More concretely, what is the relation between the Jacobians across different tokens and transformer layers?
+Q2. How do the properties of hidden representations and their relations emerge with training?
+Q3. Are any of these properties related to the generalization capabilities of LLMs?
+Figure 1: Relationship between Coupling and Benchmark Score.
+The plot illustrates the relationship between average coupling (negative mis-coupling) and benchmark score across various base LLMs (that are not fine-tuned).
+The data shows a noticeable trend: increased alignment correlates with improved benchmark scores.
+The regression fit produces an $R^{2}$ value of $0.7214$.
 1.1 Contributions
-We uncover Transformer Alignment (TA), characterized by two properties (analogous to those of RA) that consistently emerge with training in 38 openly available base and fine-tuned LLMs (Appendix A.1) including the Falcon [Almazrouei et al., 2023], Llama-3 [AI@Meta, 2024], Llama-2 [Touvron et al., 2023], MPT [Team, 2023a, b], Mistral v0.1 [Jiang et al., 2023], Gemma [Team et al., 2024], NeoX [Black et al., 2022], Neo [Black et al., 2021, Gao et al., 2020], and Pythia [Biderman et al., 2023] families.
-The properties are given by
-  • (TA1)
-    Emergence of linearity and exponentially distanced trajectories of intermediate representations in layer-wise progression.
-  • (TA2)
-    Alignment of top left and right singular vectors in the linearizations of transformer blocks through their Residual Jacobians.
-Due to the large vocabulary size relative to the embedding size in LLMs, Residual Jacobians maintain rank of at most the vocabulary size, and so the previous (RA3) immediately holds.
-Further, trained LLMs do not exhibit a consistent inverse scaling of top singular values with depth, and scaling varies considerably between models, demonstrating a lack of (RA4).
-In contrast to ResNets, the layer-wise progression of transformer representations is better described with an exponential evolution, rather than an equispaced linear progression.
-Additionally, we explore the relationship between (TA2) and generalization, and uncover that increased alignment of Residual Jacobian singular vectors correlates with improved Open LLM leaderboard [Beeching et al., 2023] benchmark score (Figure 9), raising valuable insight to the mathematical properties of LLM transformers and their connections to generalization.
+We investigate the motivating questions across several openly-available LLMs, most having over 1B parameters, trained by 6 independent organizations, with varying training methods and data (Appendix A.1).
+Through our experiments, we identify consistent patterns that define the transformer block coupling phenomenon.
+1.
+Coupling.
+The singular vectors of the Jacobians of transformer blocks couple across depth (Figures LABEL:fig:paris_all_8to16, LABEL:fig:paris_all_8to16_appendix) and tokens (Figures LABEL:fig:coupling_inoutsame, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout, LABEL:fig:coupling_inoutsame_appendix, LABEL:fig:coupling_fixed_input_appendix, LABEL:fig:coupling_fixed_output_appendix).
+Further, coupling across Jacobians emerges with training (Figures 2, LABEL:fig:paris_all_8to16, LABEL:fig:coupling_inoutsame, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout).
+2.
+Generalization.
+The strength of coupling is correlated with benchmark performance on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) (Figures 1, LABEL:fig:alignment_vary_k).
+Additionally, coupling is more strongly correlated with generalization than parameter budget, model depth, and token embedding dimension (Figure LABEL:fig:correlation_hyperparams).
+3.
+Regularity.
+Linearity in hidden trajectories emerges with training (Figures LABEL:figs:lss-training, LABEL:figs:pca, 13, 15, 23, 24).
+Exponential growth occurs in contiguous token representations as a function of depth (Figures 14, 22, 23, LABEL:figs:expo-training).
+We provide a new perspective on token embedding interactions within LLMs by examining transformers layers through their Jacobians.
+Our results display the effect of training on transformer blocks, and suggest potential approaches for promoting generalization in LLMs.
+Figure 2: Emergence of Singular Vector Coupling with Training.
+Normalized coupling (Section 3.1) evaluated for $J_{i,j}^{d_{1}},J_{i,j}^{d_{2}}$, with $|d_{1}-d_{2}|\leq 5$.
+Measurements are taken for checkpoints of Pythia 12B, Pythia 6.9B, Pythia 1B (Biderman et al., 2023) at training steps $\{1,16,128,13\text{k},23\text{k},33\text{k},\ldots,143\text{k}\}$.
+Coupling is low at early training iterations, and increases with training, with Pythia 12B and 6.9B displaying greater coupling than Pythia 1B in final training steps.
 2 Background on Large Language Models
+We describe LLMs as a deep composition of functions that iteratively transform token embeddings.
 In the input layer, $l=0$, textual prompts undergo tokenization and are combined with positional encodings to create an initial high-dimensional embedding, denoted by $x_{i}^{0}\in\mathbb{R}^{d_{\text{model}}}$ for the $i^{\text{th}}$ token.
-When these embeddings are stacked together, they form a matrix:
-$X^{0}=(x_{1}^{0},x_{2}^{0},\ldots,x_{n}^{0})\in\mathbb{R}^{d_{\text{model}}\times n}.$ (1)
+When these embeddings are stacked, they form a matrix:
+$X^{0}=(x_{1}^{0},x_{2}^{0},\ldots,x_{n}^{0})\in\mathbb{R}^{n\times d_{\text{model}}}.$ (1)
 The embeddings then pass through $L$ transformer blocks:
 $X^{0}\xrightarrow{F_{\text{block}}^{1}}X^{1}\xrightarrow{F_{\text{block}}^{2}}\cdots X^{L-1}\xrightarrow{F_{\text{block}}^{L}}X^{L}.$ (2)
-Here, $X^{l}=f_{\text{block}}^{l}(X^{l-1})$ denotes the embeddings after the $l^{\text{th}}$ block, consisting of causal multi-headed attention (MHA), a feed-forward network (FFN), and normalization layers (LN) with residual connections:
+$X^{l}=F_{\text{block}}^{l}(X^{l-1})$ denotes the embeddings after the $l^{\text{th}}$ block, consisting of causal multi-headed attention (MHA), a feed-forward network (FFN), and normalization layers (LN) with residual connections:
 $h^{l+1}(X^{l})$ $=\text{MHA}(\text{LN}(X^{l}))$ (3)
 $g^{l+1}(X^{l})$ $=\text{LN}(X^{l}+h^{l+1}(X^{l}))$ (4)
-$f_{\text{block}}^{l+1}(X^{l})$ $=X^{l}+h^{l+1}(X^{l})+\text{FFN}(g^{l+1}(X^{l})),$ (5)
-where the MHA, LN, FFN are implicitly indexed by layer. In the final representation, an additional layer normalization is applied:
+$f^{l+1}(X^{l})$ $=h^{l+1}(X^{l})+\text{FFN}(g^{l+1}(X^{l}))$ (5)
+$F_{\text{block}}^{l+1}(X^{l})$ $=X^{l}+f^{l+1}(X^{l}),$ (6)
+where the MHA, LN, FFN are implicitly indexed by layer.
+Among many models (Appendix 1), an additional rotary positional embedding (RoPE, Su et al.
+(2023)) is applied in the MHA layer.
+In the final representation, typically an additional layer normalization is applied:
 $F_{\text{block}}^{L}(X^{L-1})$ $=\text{LN}(X^{L-1}+h^{L}(X^{L-1})+\text{FFN}(g^{L}(X^{L-1}))).$ (7)
@@
 3 Methods
-3.1 Suite of Large Language Models
-Our study evaluates a total of 38 LLMs (24 base LLMs and 14 fine-tuned, see Appendix A.1) that were independently trained by various individuals and organizations.
-These models, provided through HuggingFace [Wolf et al., 2020], vary in terms of parameter budgets, number of layers, hidden dimensions, and training tokens.
-A summary of the models under consideration is presented in Table 1 of Appendix A.1.
-Measurements are made for models with varying parameter sizes with GPUs satisfying appropriate memory requirements, further detailed in Appendix A.3.
-3.2 Prompt Data
-We evaluate these LLMs using prompts of varying length, ambiguity, and context, sourced from the test set of ARC [Clark et al., 2018], GSM8K [Cobbe et al., 2021], HellaSwag [Zellers et al., 2019], MMLU [Hendrycks et al., 2021], Truthful QA [Lin et al., 2022], and Winogrande [Sakaguchi et al., 2019].
-These datasets set the performance benchmarks on the HuggingFace Open LLM Leaderboard [Beeching et al., 2023] since they encompass a diverse set of language tasks.
-Post-prompting, we assess the models using several metrics, detailed in the following subsections.
-3.3 Alignment of Singular Vectors of Residual Jacobians
-To demonstrate the presence of (TA2), we examine the properties of the transformer blocks and the relationships between them.
-This is done by analyzing the linearizations of the blocks given by their Residual Jacobian matrices
-$J_{l}^{i}=\frac{\partial}{\partial x_{i}^{l-1}}\left(h^{l}(X^{l-1})+\text{FFN\textsubscript{$l$}}(g^{l}(X^{l-1}))\right)_{i},$ (7)
-defined for each $l=1,\ldots,L$, and $i=1,\ldots,n$.
-Note that this is the Jacobian matrix for each transformer block without the contribution from the skip connection from the input of the block, analogous to the quantities measured by Li and Papyan [2023].
-The singular value decompositions of these $J_{l}^{i}$ are computed, i.e., $J_{l}=U_{l}S_{l}V_{l}^{\top}$ (with superscript $i$, indicating the token, omitted for clarity), where $U_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ and $V_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ are the matrices of left and right singular vectors respectively, and $S_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ is the singular value matrix.
-The matrices $A_{i,j,K}\coloneqq U_{j,K}^{\top}J_{i}V_{j,K}$ are plotted over all pairs of depths $i,j\in\{1,\ldots,L\}$, where $U_{j,K}$ and $V_{j,K}$ are the sub-matrices with columns that are the top-$K$ left and right singular vectors of $J_{j}$, respectively.
-To quantify the misalignment of singular vectors of the Residual Jacobians at depths $i$ and $j$, we measure how far the $A_{i,j,K}$ are from being diagonal.
-More precisely, we define the misalignment of $A_{i,j,K}$ as
-$m(A_{i,j,K})=\lVert A_{i,j,K}-\text{Diag}(A_{i,j,K})\rVert_{F},$ (8)
-where $\text{Diag}(A_{i,j,K})$ is the matrix $A_{i,j,K}$ with all non-diagonal entries replaced by zero and $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm.
-3.4 Layer-Wise Exponential Growth
-To demonstrate the presence of (TA1), we measure the expodistance of the hidden trajectories, defined below.
-In this context, each layer in the model corresponds to a unit time step of a dynamical system, so we denote $\dot{x}_{i}(t)=x_{i}(t+1)-x_{i}(t)$ the time derivative of the $i^{\text{th}}$ token.
-Assuming exponential growth of the embedding norms as they flow through the hidden layers, we can write,
-$\|\dot{x}_{i}(t)\|$ $\approx e^{\alpha t}\|\dot{x}_{i}(0)\|$ (9)
-for some fixed $\alpha\in\mathbb{R}$. We quantify the validity of the approximation in Equation (9) by measuring the coefficient of variation of $\alpha$ for each trajectory. Namely, following (9),
-$\alpha_{t}^{i}\approx\frac{1}{t}\ln\left(\frac{\|\dot{x}_{i}(t)\|}{\|\dot{x}_{i}(0)\|}\right)$ (10)
-for each time (layer) $t$ and token $i$.
-Under exponential growth, it is expected that $\alpha_{t}^{i}$ is independent of time.
-We then denote the expodistance (ED) of the trajectory of the $i^{th}$ token of a given sequence by
-$ED_{i}=\frac{\text{Var}_{t}\alpha_{t}^{i}}{(\text{Avg}_{t}\alpha_{t}^{i})^{2}}$ (11)
-This measurement is motivated by the parametrization discussed in Appendix A.2 and serves as a method to test the validity of the linearization presented in Equation (15).
+3.1 Coupling of Singular Vectors of Jacobians
+Jacobians. Coupling is investigated through analyzing the linearizations of transformer blocks given by their Jacobians
+$J_{t_{1}t_{2}}^{l}=\frac{\partial}{\partial x_{t_{1}}^{l-1}}\left(f^{l}(X^{l-1})\right)_{t_{2}},$ (8)
+defined for each layer $l\in\{1,\ldots,L\}$, and pair of tokens $t_{1},t_{2}\in\{1,\ldots,n\}$.
+Note that this is the Jacobian matrix for each transformer block without the contribution from the skip connection from the input of the block, similar to the quantity measured by Li & Papyan (2023), but extended for cross-token interactions.
+Also, due to the causal structure of the representations, $J_{t_{1}t_{2}}^{l}=0$ if $t_{1}>t_{2}$, so we restrict our attention to the case $t_{1}\leq t_{2}$.
+Singular value decomposition.
+The SVD of the Jacobians $J_{t_{1}t_{2}}^{l}$ are computed, i.e., $J_{t_{1}t_{2}}^{l}=U_{l}S_{l}V_{l}^{\top}$ (with superscripts $t_{1},t_{2}$, indicating the tokens, omitted for clarity), where $U_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ and $V_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ are matrices of left and right singular vectors respectively, and $S_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ is the singular value matrix.
+To measure coupling between Jacobians $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}}$, define $A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime},K}\coloneq U_{l^{\prime},K}^{\top}J_{t_{1}t_{2}}^{l}V_{l^{\prime},K}$ for $l,l^{\prime}\in\{1,\ldots,L\}$ and $t_{1},t_{2},t_{1}^{\prime},t_{2}^{\prime}\in\{1,\ldots,n\}$, where $U_{l^{\prime},K}$ and $V_{l^{\prime},K}$ are the sub-matrices with columns that are the top-$K$ left and right singular vectors of $J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}}$, respectively.
+Figure 3: Transformer Block Coupling.
+Displayed is a visualization of the various types of transformer block coupling, along with brief instructions on computing both the Jacobians $J$ and coupling matrices $A$.
+Depth-wise Coupling.
+To analyze coupling across transformer blocks, we fix $t$ (usually as the last token), and check alignment between $J_{tt}^{l}$ and $J_{tt}^{l^{\prime}}$ through $A_{t}^{ll^{\prime},K}$ (we write the single subscript $t$ for clarity) across layers $l,l^{\prime}\in\{1,\ldots,L\}$.
+Token-wise Coupling.
+To study the coupling across tokens, we start by fixing two layers $l,l^{\prime}\in\{1,\ldots,L\}$, and analyze coupling across different cases of the $t_{1},t_{2},t_{1}^{\prime},t_{2}^{\prime}$.
+We start with the case where the input and output tokens are the same and compare $J_{tt}^{l}$ and $J_{t^{\prime}t^{\prime}}^{l^{\prime}}$ across $t,t^{\prime}\in\{1,\ldots,n\}$, which is the coupling across tokens for a token’s effect on its own trajectory.
+We also consider the context tokens’ impact on a trajectory by including coupling across cases with $t_{1}<t_{2}$.
+More specifically, we measure coupling between $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}t_{2}^{\prime}}^{l^{\prime}}$ across $t_{2},t_{2}^{\prime}\geq t_{1}$ (fixing the input token to be the same) and also between $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}^{\prime}t_{2}}^{l^{\prime}}$ across $t_{1},t_{1}^{\prime}\leq t_{2}$ (fixing the output token to be the same).
+Coupling measurement.
+To quantify the coupling of singular vectors of the Jacobians, we measure how far the $A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime},K}$ are from being diagonal.
+More precisely, we quantify the mis-coupling of $A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime},K}$ using $m:\mathbb{R}^{K\times K}\to\mathbb{R}$ defined by
+$m(A)=\lVert A-\text{Diag}(A)\rVert_{F},$ (9)
+where $\text{Diag}(A)$ is the matrix $A$ with all non-diagonal entries replaced by zero and $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm.
+For normalized comparison between models across training in Figure 2, we normalize $A$ by their norms
+$\tilde{m}(A)=\frac{\lVert A-\text{Diag}(A)\rVert_{F}}{\lVert A\rVert_{F}}$ (10)
 3.2 Linearity of Trajectories
-Linearity in intermediate embeddings (TA1) is quantified with the line-shape score (LSS), defined by Gai and Zhang [2021] as
+Linearity in intermediate embeddings is quantified with the line-shape score (LSS), defined by Gai & Zhang (2021) as
 $\text{LSS}_{i}^{0,\ldots,L}=\frac{L}{\left|\left|\tilde{x}^{L}_{i}-\tilde{x}^{0}_{i}\right|\right|_{2}},$ (11)
@@
 Note that $\text{LSS}\geq 1$, with $\text{LSS}=1$ if and only if the intermediate representations $x_{i}^{0},\ldots,x_{i}^{L}$ form a co-linear trajectory.
-3.6 Visualization of Trajectories with PCA
+3.3 Layer-Wise Exponential Growth
+We measure the presence of exponential spacing (expodistance) of the hidden trajectories. Assuming exponential growth of the embedding norms as they flow through the hidden layers, we can estimate,
+$\|x_{i}^{l}\|$ $\approx e^{\alpha l}\|x_{i}^{0}\|=e^{\alpha}\|x_{i}^{l-1}\|$ (13)
+for some fixed $\alpha\in\mathbb{R}$ over all layers $l=1,\ldots L$.
+We quantify the validity of the representation in Equation 13 by measuring the coefficient of variation of $\alpha_{i}^{l}$, given by
+$\alpha_{i}^{l}\approx\ln\left(\frac{\|x_{i}^{l}\|}{\|x_{i}^{l-1}\|}\right),$ (14)
+for each layer $l$ and token $i$.
+Under exponential growth, it is expected that $\alpha_{i}^{l}$ is independent of depth.
+We therefore denote the expodistance (ED) of the trajectory of the $i^{th}$ token of a given sequence by
+$\text{ED}_{i}=\frac{\text{Var}_{l}\alpha_{i}^{l}}{(\text{Avg}_{l}\alpha_{i}^{l})^{2}}.$ (15)
+This measurement is motivated by the parametrization discussed in Appendix A.4, as well as empirical evidence in Figure LABEL:figs:norm-expo-traininga, and serves as a method to test the validity of the linearization presented in Equation 17.
+4 Evaluation
+4.1 Suite of Large Language Models
+Our study evaluates a total of 38 LLMs (24 base LLMs and 14 fine-tuned, see Appendix A.1) that were independently trained by various individuals and organizations.
+These models, provided through HuggingFace (Wolf et al., 2020), vary in terms of parameter budgets, number of layers, hidden dimensions, and training tokens.
+Moreover, we analyze the dynamics of each measurement throughout training by deploying the Pythia Scaling Suite (Biderman et al., 2023).
+A summary of the models under consideration is presented in Table 1 of Appendix A.1 and further details in Appendix A.5.
+4.2 Prompt Data
+We evaluate these LLMs using prompts of varying length, ambiguity, and context, sourced from the test set of ARC (Clark et al., 2018), GSM8K (Cobbe et al., 2021), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), Truthful QA (Lin et al., 2022), and Winogrande (Sakaguchi et al., 2019).
+This data sets the performance benchmarks on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) and provide a representative evaluation of performance on many language tasks.
+5 Results
+5.1 Coupling of Jacobians Across Depth
+In trained LLMs, we observe coupling of the top singular vectors of the Jacobians across depth (Figure LABEL:fig:paris_all_8to16 bottom row), evident in the low non-diagonal values with a visible diagonal present in the matrix subplots.
+This is consistently observed across various LLMs considered.
+On the other hand, in untrained models (Figure LABEL:fig:paris_all_8to16 top row), there is no coupling of Jacobians across different depths.
+There is coupling along the diagonal, however, because each Jacobian is trivially diagonalized by its own singular vectors.
+This, in addition to Figure 2, suggests that coupling across depth emerges through training.
+5.2 Coupling of Jacobians Across Tokens
+We also analyze the coupling of singular vectors of the Jacobians across tokens.
+First, we observe coupling between Jacobians where the input and output tokens are the same, that is, between $J_{l}^{tt}$ and $J_{l^{\prime}}^{t^{\prime}t^{\prime}}$ across $t,t^{\prime}\in\{1,\ldots,n\}$ (Figure LABEL:fig:coupling_inoutsame).
+This suggests that a token’s interaction on its own trajectory is coupled with that of other tokens.
+We also consider the context tokens’ interaction on a trajectory when fixing the input token $t_{1}$ (Figure LABEL:fig:coupling_fixedin), through $J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}t_{2}^{\prime}}$ across $t_{2},t_{2}^{\prime}\geq t_{1}$, and when fixing the output token $t_{2}$ (Figure LABEL:fig:coupling_fixedout), with $J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}^{\prime}t_{2}}$ across $t_{1},t_{1}^{\prime}\leq t_{2}$.
+Although generally present, the coupling in these two cases are less consistent, with variation in the strength of coupling between different token pairs.
+Again in untrained models (Figure LABEL:fig:coupling_inoutsame top row), there is a lack coupling of Jacobians across tokens.
+5.3 Emergence of Coupling with Training
+Coupling emerges through training for the evaluated LLMs, including coupling across depth (Figure LABEL:fig:paris_all_8to16 ) and across tokens (Figures LABEL:fig:coupling_inoutsame, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout).
+Further, we evaluate layer-wise coupling at intermediate training checkpoints of Pythia 12B, 6.9B, and 1B (Biderman et al., 2023) (Figure 2), and observe that coupling is generally low at initialization and increases significantly during early training steps.
+5.4 Correlation with Generalization
+For each LLM, we measure the average coupling across depth(which we define to be negative mis-coupling) across prompts in the 6 evaluation datasets (Section 4.2), where for each prompt, $m(A_{ll^{\prime},K}^{n})$ is averaged over layers $l,l^{\prime}\in\{1,\ldots,L\}$.
+We plot the coupling values against the benchmark scores across several LLMs (Figure 1).
+Our results reveal a positive correlation between coupling and performance benchmark scores, and is more significant than the relationship between other significant model hyperparameters (Figure LABEL:fig:correlation_hyperparams).
+This observation suggests a compelling relationship between stronger coupling of singular vectors of Jacobians $J_{l}^{t_{1}t_{2}}$ and improved generalization.
+5.5 Regularity in Hidden Trajectories
+We identify a considerable degree of linearity in the hidden trajectories of all featured LLMs.
+The LSS of the trajectories has an average value $4.25$ in trained models, while taking average values of $6.54$ at initialization, and is supported by the low variation across benchmark prompts (Figure 13).
+Linearity increases with training at varying depths of Pythia 12B (Figure LABEL:figs:lss-training).
+Linear and expansive behavior of the representations is demonstrated in Llama-3 70B, Mistral 30B, and NeoX 20B through low dimensional projections of embedding trajectories (Figure LABEL:figs:pca).
+Figure 13: Linearity of Trajectories.
+The figure depicts the line-shape score (LSS) of embedding trajectories, as discussed in Section 5.5), computed on 1,200 prompts of the HuggingFace Open LLM Leaderboard (Section 4.2) for a variety of trained (black) and randomly initialized (red) LLMs (Appendix A.1).
+Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model.
+Models are sorted by number of parameters.
+5.6 Layer-Wise Exponential Growth
+Across all LLMs considered, the hidden trajectories exhibit exponential growth that emerges with training (Figure 14).
+The majority of the models exhibit a low coefficient of variation across prompts, showing the robustness of this property across a variety of tasks.
+The expodistance measurements show little dependence on model parameters, and take on consistently low values across all featured LLMs (Figure 14).
+In contrast, measurements at initialization show equally (rather than exponentially) distanced trajectories, as reflected by the low coefficients of variation in the norms of their layer-wise differences (Figure 22).
+Under certain assumptions, exponential spacing is justified in Section 6.2.
+Figure 14: Coefficient of variation of layer-wise exponent.
+The figure presents the coefficient of variation of the layer-wise exponent (Section 3.3) computed on 1,200 prompts of the HuggingFace Open LLM Leaderboard (Section 4.2) for trained LLMs (Appendix A.1).
+Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model.
+LLMs are sorted by increasing benchmark performance, while the vertical axis has a base 10 logarithmic scale.
+6 Discussion
+6.1 Correlation with Generalization
+As discussed in Section 5.4, stronger coupling across Jacobian singular vectors is related to improved generalization (Figure 1).
+This relationship may assist in evaluating the training of transformers.
+In particular, this finding suggests that the development of training methods to amplify coupling across transformer blocks may lead to favourable model performance.
+6.2 Emergence of Exponential Spacing with Coupling
+The emergence of exponential spacing in many LLMs can be viewed as a result of coupling under some assumptions on the homogeneity of the singular values across depth.
+Strong coupling suggests that the linearizations of hidden trajectories satisfy a linear difference equation (see Appendix A.4), given by
+$x^{l+1}-x^{l}=Ax^{l}$
+Considering the singular value decomposition of $A$ in Equation 18 as $A=U\Sigma V^{T}$, and assuming that there is coupling ($U=V$), the difference equation can be diagonalized in the same basis across depth.
+Further assuming that the top singular values of the Jacobian blocks display low variation across depth, (Figure LABEL:fig:svalues_trained), we expect
+$||x^{l}||\approx e^{\sigma l}\,||x^{0}||,$ (16)
+where $\sigma$ is the top singular value.
+Since we assume $\sigma$ is independent of depth, the above norm is expected to be exponential in $l$, and offers an explanation to the measurements in Figure 14.
+7 Related Work
+Residual Networks.
+ResNets (He et al., 2016) have been viewed as an ensemble of shallow networks (Veit et al., 2016), with studies delving into the scaling behaviour of their trained weights (Cohen et al., 2021).
+The linearization of residual blocks by their Residual Jacobians was first explored by Rothauge et al.
+(2019), who examined Residual Jacobians and their spectra in the context of stability analysis, and later by Li & Papyan (2023) who discovered Residual Alignment.
+Coupling of Jacobian singular vectors in LLM transformers extends previous results for classifier Resnets (Li & Papyan, 2023).
+We show coupling in $J_{l}^{t_{1}t_{2}}$ across various tokens, which is specific to tokenization in transformers, in addition to demonstrating coupling of $J_{l}^{t_{1}t_{2}}$ across $l$, which was also identified analogously in ResNets (Li & Papyan, 2023).
+Our measurements quantify the emergence with training of linearity and spacing, including at different depths, which was previously observed mostly qualitatively.
+Further comparison is included in Appendix A.3.
+Neural ODE.
+Introduced by (Chen et al., 2018), Neural ODE view ResNets as a discretized dynamical process, with past work (Sander et al., 2022) showing the convergence of Residual Networks to a system of linear ODE.
+(Zhong et al., 2022; Li et al., 2021) studied neural ODE specific perspective to transformers and other LLM architectures.
+In our work, the presence of residual connections and emergence of Jacobian Alignment in transformers suggests that a discretization of a simple ODE emerges in LLMs.
+Information Processing in LLMs.
+Previously, Katz & Belinkov (2023); Bietti et al.
+(2023) have described memory and semantic flow in intermediate states of transformers while van Aken et al.
+(2019) have examined BERT hidden trajectories.
+Investigation of LLMs for reasoning tasks and interpretability remains an important research focus (Huang & Chang, 2023).
+In-context learning.
+Previous works have observed that LLMs are able to perform in context learning (von Oswald et al., 2023; Bai et al., 2024; Ahn et al., 2023; Akyürek et al., 2023; Xie et al., 2021; Hahn & Goyal, 2023; Xing et al., 2024): the model is capable of performing a particular machine learning task through a few examples provided in a single prompt or context.
+Their results consider a single layer of trained self-attention as implementing a single step of a gradient-descent-like learning rules across transformers depth, minimizing the MSE of a linear model
+$x_{\text{final}}=\min_{x}||Ax-b||^{2}$
+A gradient-descent learning rule would cause depth-wise updates to representations of the form
+$x_{t+1}=x_{t}+\epsilon A^{T}(Ax_{t}-b)=(I+\epsilon A^{T}A)x_{t}-{\epsilon}A^{T}b$
+We should expect similarity in the matrices $I+{\epsilon}A^{T}A$ across depth, which is suggested by coupling.
+Hidden Representation Dynamics.
+Previous studies (Gai & Zhang, 2021; Haber & Ruthotto, 2017; Ee, 2017), have focused on interpreting deep neural networks from a dynamical systems perspective.
+Throughout training, numerical experiments have shown that both plain networks and ResNets are aptly approximated by a geodesic curve in Wasserstein space (Gai & Zhang, 2021).
+Other interpretations involve partitioning activation space into basins of attraction, with fixed points representing discrete thoughts (Nam et al., 2023).
+For further related works, see (Geshkovski et al., 2023b; a; Tarzanagh et al., 2023; Valeriani et al., 2023).
+Structure in Hidden Representations.
+Emergence of regularity in last-layer representations during training was noticed in Neural Collapse (Papyan et al., 2020).
+Subsequent works examined structure of hidden representations and presented theoretical results describing the emergence of Neural Collapse (Wang et al., 2024a; Parker et al., 2023; Zangrando et al., 2024; Garrod & Keating, 2024; Wang et al., 2024b; Hoyt & Owen, 2021; Arous et al., 2023; Zarka et al., 2021; Ben-Shaul & Dekel, 2022; Papyan, 2020; Súkeník et al., 2023).
+Within LLMs, recent works have identified uniform structure in token means (Wu & Papyan, 2024), and low dimensional structure across hidden trajectories (Sarfati et al., 2024).
+Our work studies the representations of all layers in LLMs in a local sense through the interactions of tokens through their Jacobians.
+8 Limitations
+A limitation of the scope of our analysis is that due to the significant compute resources required to train LLMs of the size studied in this paper, our exploration is restricted to pretrained LLMs and those fine-tuned on them.
+Several of these models are trained independently of each other, without carefully controlling for different experimental parameters.
+Consequently, direct comparisons between models become challenging, and so determining the cause of varying levels of regularity between LLMs becomes difficult.
+Nonetheless, this makes the consistent emergence of these properties interesting for further study.
+9 Conclusion
+Our primary goal was to contribute to the understanding of the mechanics underlying transformer architectures through an analysis of the trajectories of token embeddings and their interactions.
+Our research builds on the understanding of transformer architectures by revealing the coupling, across depth and token, of singular vectors in the Jacobians of transformer blocks for multiple LLMs trained by various organizations.
+We additionally observe and quantify the emergence of linearity and exponential spacing with training in the embedding trajectories of LLMs.
+Notably, we establish a correlation between the strength of this coupling and benchmark performance on the HuggingFace Open LLM Leaderboard, highlighting the significance of transformer block coupling for generalization.
+These findings open avenues for future research, encouraging deeper exploration into the connections between regularity of hidden representations, model specifications, and generalization.
+Appendix A Appendix and Supplementary material
+A.1 Suite of Large Language Models and Prompt Data
+We evaluate 38 total LLMs (24 base LLMs and 14 fine-tuned) on 6 datasets from the HuggingFace Open LLM leaderboard (Beeching et al., 2023).
+Base models.
+Falcon (40B, 7B) (Almazrouei et al., 2023), Llama-3 (70B, 8B) (AI@Meta, 2024), Llama-2 (70B, 13B, 7B) (Touvron et al., 2023), MPT (30B, 7B) (Team, 2023a; b), Mistral v0.1 (Jiang et al., 2023), Gemma (7B, 2B) (Team et al., 2024), Gemma 1.1 (7B, 2B), NeoX (20B) (Black et al., 2022), Neo (2.7B) (Black et al., 2021; Gao et al., 2020), Pythia (6.9B).(Biderman et al., 2023), and GPT-2 (1.5B, 774M, 355M, 117M) (Radford et al., 2019).
+Fine-tuned models.
+CodeLlama (34B, 13B, 7B) (Rozière et al., 2024), CodeLlama Instruct (34B, 13B, 7B) (Rozière et al., 2024), Mistral-v0.1 Instruct (7.3B) (Jiang et al., 2023), Mistral-v0.2 Instruct (Jiang et al., 2023), CodeGemma (Team et al., 2024).
+Datasets.
+ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), Truthful QA (Lin et al., 2022), WinoGrande (Sakaguchi et al., 2019).
+Table 1: LLMs featured in the experiments throughout paper.
+Included in the table is the parameter budget of each model, the embedding dimension, the number of training tokens, and the Open LLM leaderboard (Beeching et al., 2023) benchmark score.
+Model Param. Layers ($L$) Dim. ($d_{\text{model}}$) Tokens Score Pos.Encoding
+Llama-3 70 B 80 8192 15 T RoPE
+8 B 32 4096 15 T 62.35 RoPE
+Llama-2 70 B 80 8192 2 T 66.05 RoPE
+13 B 40 5120 2 T 55.69 RoPE
+7 B 32 4096 2 T 50.97 RoPE
+Codellama 34 B 48 8192 2 T 55.33 RoPE
+13 B 40 5120 2 T 45.82 RoPE
+7 B 32 4096 2 T 39.81 RoPE
+Codellama (IT) 34 B 48 8192 2 T 43.0 RoPE
+13 B 40 5120 2 T 37.52 RoPE
+7 B 32 4096 2 T 40.05 RoPE
+Orca 13 B 40 5120 2 T 58.64 RoPE
+7 B 32 4096 2 T 54.55 RoPE
+Falcon 40 B 60 8192 1 T 58.07 RoPE
+7 B 32 4544 1.5 T 44.17 RoPE
+Falcon (IT) 40 B 60 8192 1 T 43.26 RoPE
+7 B 32 4544 1.5 T RoPE
+MPT 30 B 48 7168 1 T 66.98 ALiBi
+7 B 32 4096 1 T 56.83 ALiBi
+Phi 2 B 32 2560 1.4 T 61.33 RoPE
+1.5 B 24 2048 150 B 47.69 RoPE
+1 B 24 2048 54 B RoPE
+Mistral-v0.1 7.3 B 32 4096 60.97 ALiBi
+Mistral-v0.1 (IT) 7.3 B 32 4096 54.96 ALiBi
+Mistral-v0.2 (IT) 7.3 B 32 4096 65.71 ALiBi
+Gemma 7 B 28 3072 6 T 64.29 RoPE
+2 B 18 2048 6 T 42.75 RoPE
+Gemma-1.1 7 B 28 3072 6 T 30.0 RoPE
+2 B 18 2048 6 T 60.09 RoPE
+Codegemma 7 B 28 3072 6 T 56.73 RoPE
+2 B 18 2048 6 T 32.19 RoPE
+Neo 20 B 44 6144 41.69 RoPE
+2.7 B 32 2560 0.42 36.20 Sine
+Pythia 12 B 36 5120 0.3 T 58.9 RoPE
+6.9 B 32 4096 0.3 T 39.30 RoPE
+2.8 B 32 2560 0.3 T RoPE
+1.4 B 24 2048 0.3 T RoPE
+1 B 16 2048 0.3 T RoPE
+GPT-2 1.5 B 48 1600 34.12 Sine
+774 M 36 1280 32.07 Sine
+355 M 24 1024 29.87 Sine
+117 M 12 768 28.53 Sine
+A.2 Additional Metrics
+A.2.1 Visualization of Trajectories with PCA
 Each token, with initial embedding $x_{i}^{0}$, forms a trajectory $x_{i}^{0},x_{i}^{1},\ldots,x_{i}^{L}$ as it passes through the $L$ transformer blocks.
@@
 The projected embeddings, $\text{PC}_{L}(x_{i}^{0}),\,\text{PC}_{L}(x_{i}^{1}),\,\ldots,\,\text{PC}_{L}(x_{i}^{L})$, are plotted for each of the $i=1,\ldots,n$ trajectories.
-4 Results
-4.1 Alignment of Singular Vectors of Residual Jacobians
-In trained LLMs, we observe (TA2), that is, alignment of the top singular vectors of the Jacobians $J_{l}$ across depth (Figure LABEL:fig:paris_all_8to16 bottom row), evident in the distinct diagonal lines present in the matrix subplots.
-This phenomenon is consistently observed across LLMs, similar to (RA2), which was observed by Li and Papyan [2023] in ResNets.
-On the other hand, in untrained models (Figure LABEL:fig:paris_all_8to16 top row), there is no alignment between singular vectors of Residual Jacobians across different depths.
-There is alignment along the diagonal, however, because each Residual Jacobian is trivially diagonalized by its own singular vectors in $A_{i,j,K}$.
-This suggests that (TA2), the alignment of singular vectors of Residual Jacobians, emerges through training.
-To understand this phenomenon more precisely, for each LLM, we measure the average alignment (which we define to be negative misalignment) across prompts in the 6 evaluation datasets (Section 3.2), where for each prompt, the misalignment $m(A_{i,j,K})$ is averaged over layers $i,j=1,\ldots,L$.
-We plot the negative misalignment values against the benchmark scores across several LLMs (Figure 9).
-Our results reveal a consistent pattern: an increase in alignment corresponds to an increase in benchmark scores.
-This observation suggests a compelling relationship between stronger alignment of singular vectors of Residual Jacobians and improved generalization performance.
-Figure 9: Relationship between Alignment and Benchmark Score.
-The plot illustrates the relationship between average alignment (negative misalignment) and benchmark score across various LLMs.
-The left subplot includes models fine-tuned for coding and instruction, while the right excludes these models.
-We make this distinction as these fine-tuned models are for code and instruction, rather than being fine-tuned on the benchmark tasks.
-The data shoes a noticeable trend: increased alignment correlates with elevated benchmark scores.
-In particular, when excluding the code and instruction models, the linear regression gives an $R^{2}$ value of $0.7214$.
-This suggests that stronger singular vector alignment may correspond with enhanced generalization performance.
-4.2 Linearity of Trajectories
-We observe (TA1) in each LLM, evident in the LSS of layer-wise trajectories at initialization and after training evaluated over benchmark prompts (Figure 10), in agreement with (RA1) in ResNets.
-Linear and expansive dynamical behavior is demonstrated in Llama-3 70B, Mistral 30B, and NeoX 20B through low dimensional projections of embedding trajectories (Figure LABEL:figs:pca).
-The LSS of the trajectories has an average value $4.25$ across trained models, while taking average values of $6.54$ at initialization, and is evidenced by the low variation across benchmark prompts (Figure 10).
-Figure 10: (TA1): Linearity of layer-wise trajectories.
-The figure depicts the line-shape score (LSS) of embedding trajectories, as discussed in Section 4.2), computed on 1,200 prompts of the HuggingFace Open LLM Leaderboard (Section 3.2) for a variety of trained (black) and randomly initialized (red) LLMs (Appendix A.1).
-Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model.
-Models are sorted by increasing benchmark performance.
-Across every model, training increases linearity (lower LSS) of embedding trajectories when compared to that of untrained models (greater LSS).
-After training, LSS appears largely independent of the parameter budget, embedding dimension, and the HuggingFace OpenLLM benchmark score.
-At initialization, the non-linearity of the trajectories grows with model size, embedding dimension, and number of layers (Figure 10).
-Moreover, the discrepancy between trained and untrained models increases with the number of parameters, hidden layers, and embedding dimension.
-4.3 Layer-wise Growth
-Across all LLMs considered, the hidden trajectories exhibit distinctly exponential growth that emerges with training (Figure 11), displaying the second property of (TA1).
-This property, which arises as a result of Jacobian alignment, is distinct from (RA1) in the sense that distances between consecutive layers grow exponentially instead of linearly with depth, and highlights a key difference between the trajectories in ResNet classifiers and transformers.
-Most of the LLMs considered exhibit a low coefficient of variation across prompts, showing the robustness of this property across a variety of tasks.
-The expodistance measurements show little dependence on model parameters, and take on consistent and small values across the models (Figure 11).
-Figure 11: (TA1): Coefficient of variation of layer-wise exponent.
-The figure presents the coefficient of variation of the layer-wise exponent (Section 3.4) computed on 1,200 prompts of the HuggingFace Open LLM Leaderboard (Section 3.2) for trained LLMs (Appendix A.1).
-Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model.
-LLMs are sorted by increasing benchmark performance, while the vertical axis has a base 10 logarithmic scale.
-The coefficient of variation for most models is low, suggesting that the growth of the embedding norms follows an exponential pattern across different layers, characterized by a relatively constant exponent.
-At initialization, each model consistently displays equally (rather than exponentially) distanced trajectories, as reflected by the extremely low coefficients of variation in the norms of their layer-wise differences (Figure 23).
-Conversely, the trajectories of the trained models appear to be significantly less equidistant across layers, exhibiting substantially larger coefficients of variation with far less consistency across prompts.
-This behaviour, however, is expected and is discussed in more detail in Section 5.4.
-5 Discussion
-5.1 Correlation with Generalization
-As discussed in Section 4.1, stronger alignment of Residual Jacobians singular vectors is related to improved generalization (Figure 9).
-This relationship between (TA2) with LLM benchmark scores may assist in evaluating the training of transformers.
-Understanding the relationship between training methods and regularity in representations holds potential implications for the enhancement of LLMs.
-This finding suggests that development of techniques to amplify Transformer Alignment may lead to favourable model performance.
-The connection between hidden representation regularity, generalization, and LLM training, as prompted by our work, presents an intriguing research direction with important practical implications.
-5.2 Regularity and Training
-Many LLMs are based on the same transformer architecture, yet when trained by independent individuals and organizations, training directly affects the emergence of residual Jacobian alignment (Figure 9), presence of linearity (Figure 10), and extent of exponential growth (Figure 11).
-The effect of training is further evident when comparing base LLMs to their fine-tuned counterparts, since differences in measurements are a direct product of further optimization of the same weights.
-Our findings further highlight the important effect of training on the dynamical properties of hidden trajectories in transformers.
-5.3 Comparison to ResNets
+A.3 Comparison to ResNets
+Coupling.
+Our results complement and build upon those of (Li & Papyan, 2023), who have observed the coupling of singular vectors of Residual Jacobians in classification ResNets.
+We observe coupling across depth in a wide range of LLMs, where Jacobians are evaluated at the sequence of embeddings, with respect to the current token, whereas in RA Jacobians are evaluated at a single representation.
+Additionally, with transformers we may analyze coupling not only across depth but also across tokens.
+We observe coupling in LLMs across tokens in a variety of ways.
+Further, we consider and analyze the relationship between coupling and generalization.
 Linearity and Equidistance.
-Linearity in hidden trajectories, as observed in ResNets (RA1) Li and Papyan [2023], Gai and Zhang [2021], also emerges with training in LLMs (TA1).
-The mean LSS value among the evaluated LLMs is $4.24$ (Figure 10), greater than LSS measurements observed for ResNets (Gai and Zhang [2021], page 18) which range between $2.0$-$3.0$.
+Linearity in hidden trajectories, as observed in ResNets Li & Papyan (2023); Gai & Zhang (2021), also emerges with training in LLMs.
+The mean LSS value among the evaluated LLMs is $4.24$ (Figure 13), greater than LSS measurements observed for ResNets (Gai & Zhang (2021), page 18) which range between $2.0$-$3.0$ (due to varying trajectory length and hidden dimension).
 In both architectures, training induces improved linearity and regularity (Figure LABEL:figs:pca) in trajectories.
 In contrast to ResNets, trajectories are not equidistant, instead showing exponential growth between layers (Figure 14).
-A low coefficient of variation displays the presence of (TA1) on token trajectories, in analogy with the equidistant spacing of (RA1).
-In both ResNet classifiers and LLM transformers, there is an evident level of regularity in hidden representations (Figure LABEL:figs:pca).
-Residual Jacobian Singular Vector Alignment.
-The alignment of singular vectors of Residual Jacobians is present across LLMs (Figure 9), similar to how it occurs in ResNets [Li and Papyan, 2023], and does not occur at initialization.
-Rank of Residual Jacobians.
-Li and Papyan [2023] show that residual Jacobians have rank at most $C$, the number of classes.
+We quantify this spacing through a low coefficient of variation, displaying the presence of exponential growth in token trajectories.
+In both classifier ResNets and LLM transformers, there is an evident level of regularity in hidden representations (Figure LABEL:figs:pca).
+Rank of Jacobians.
+Li & Papyan (2023) show that residual Jacobians have rank at most $C$, the number of classes.
 This analogous result automatically holds for LLMs since the vocabulary size is significantly greater than the embedding dimension of the transformer blocks.
 Singular Value Scaling.
-Li and Papyan [2023] observe that top singular values of residual Jacobians scale inversely with depth.
+Li & Papyan (2023) observe that top singular values of residual Jacobians scale inversely with depth.
 In trained LLMs, however, top singular values do not show a consistent depth scaling across models (Figure LABEL:fig:svalues_trained), notably differing from classification ResNets.
 In addition, the distribution of singular values at each layer varies significantly between models.
-Interestingly, singular value scaling is actually more present in untrained transformers (FigureLABEL:fig:svalues_untrained).
-5.4 Emergence of Exponential Growth
-Trained ResNets distinctly feature equidistance in hidden trajectories, while measurements in transformers display a layer-wise exponential growth in norm.
-[Li and Papyan, 2023] prove that equidistance in certain classifiers emerges under the assumptions of (RA 2-4), in particular, the authors assume that top singular values of residual Jacobians scale inversely with model depth.
-In transformers, top singular values do not show a similar scaling, possibly suggesting that (RA4) is a necessary condition for equidistance.
-Emergence of (TA1) appears to be a result of (TA2) and a relationship between the singular values, lying analogously to the emergence of (RA1) as a result of (RA 2-4).
-(TA2) implies that linearizations of hidden trajectories take a simple form after training, satisfying a linear difference equation (see Appendix A.2).
-In contrast to (RA4), the top singular values of the Jacobian blocks exhibit low variation across depth in some models (Figure LABEL:fig:svalues_trained).
-Under these assumptions, decomposing the matrix $A$ in Equation (16) as $A=U\Sigma V^{T}$, we obtain
-$|x(t)|\approx\exp(\sigma t)|x(0)|,$ (14)
-where $\sigma$ is the top singular value. Since $\sigma$ is independent of depth, the above norm will be exponential in $t$, and directly explains the pattern seen in Figure 11.
-Common in many models at initialization, the top singular values of the Jacobian matrices decay reciprocally with depth (Figure LABEL:fig:svalues_untrained).
-As a result, relatively linear trajectories appearing in the untrained models is not surprising.
-Randomly initialized Jacobian matrices with decaying singular values, coupled with Equation 17 results in the approximate exponential growth becoming linearized due to the receding impact of each layer.
-6 Related Work
-Residual Networks.
-ResNets have been viewed as an ensemble of shallow networks [Veit et al., 2016], with studies delving into the scaling behaviour of their trained weights [Cohen et al., 2021].
-The linearization of residual blocks by their Residual Jacobians was first explored by Rothauge et al.
-[2019], who examined Residual Jacobians and their spectra in the context of stability analysis, and later by Li and Papyan [2023] who discovered Residual Alignment.
-We continue this line of work by investigating Residual Jacobians in transformer architectures and uncover a similar phenomenon in pretrained LLMs.
-In-context learning.
-Previous works have observed that LLMs are able to perform in context learning [von Oswald et al., 2023, Bai et al., 2024, Ahn et al., 2023, Akyürek et al., 2023, Xie et al., 2021, Hahn and Goyal, 2023, Xing et al., 2024]: the model is capable of performing a particular machine learning task through a few examples provided in a single prompt or context.
-Subsequent theoretical works explained this phenomenon by postulating that intermediate representations of deep networks are capable of emulating classical learning algorithms like solving least squares through gradient descent, solving LASSO, and even training a shallow network.
-Our work provides empirical evidence for these works by demonstrating that the transformer blocks of LLMs are coupled in a way that would only emerge if consecutive blocks jointly try to solve a specific problem.
-Neural ODE.
-Introduced by [Chen et al., 2018], Neural ODE view ResNets as a discretized dynamical process, with past work [Sander et al., 2022] showing the convergence of Residual Networks to a system of linear ODE.
-[Zhong et al., 2022, Li et al., 2021] studied neural ODE specific perspective to transformers and other LLM architectures.
-In our work, the presence of residual connections and emergence of Jacobian Alignment in transformers suggests that a discretization of a simple ODE emerges in LLMs.
-Information Processing in LLMs.
-Previously, Katz and Belinkov [2023], Bietti et al.
-[2023] have described memory and semantic flow in intermediate states of transformers while van Aken et al.
-[2019] have examined BERT hidden trajectories.
-Investigation of LLMs for reasoning tasks and interpretability remains an important research focus [Huang and Chang, 2023].
-Hidden Representation Dynamics.
-Previous studies [Gai and Zhang, 2021, Haber and Ruthotto, 2017, Ee, 2017], have focused on interpreting deep neural networks from a dynamical systems perspective.
-In particular, the layer-wise transformations between transformer blocks are interpreted as discrete approximations to continuous curves through representation space.
-Throughout training, numerical experiments have shown that both plain networks and ResNets are aptly approximated by a geodesic curve in Wasserstein space [Gai and Zhang, 2021].
-Other interpretations involve partitioning activation space into basins of attraction, with fixed points representing discrete thoughts [Nam et al., 2023].
-For further related works, see [Geshkovski et al., 2023b, a, Tarzanagh et al., 2023, Valeriani et al., 2023].
-Structure in Hidden Representations.
-Emergence of regularity in last-layer representations during training was noticed in Neural Collapse [Papyan et al., 2020].
-Motivated by this phenomenon, subsequent works examined structure of hidden representations and presented theoretical results describing the emergence of Neural Collapse [Wang et al., 2024a, Parker et al., 2023, Zangrando et al., 2024, Garrod and Keating, 2024, Wang et al., 2024b, Hoyt and Owen, 2021, Arous et al., 2023, Zarka et al., 2021, Ben-Shaul and Dekel, 2022, Papyan, 2020, Súkeník et al., 2023].
-Our work studies the representations of all layers in LLMs in a local sense through individual trajectories, and uncovers regularity in the representations in LLMs.
-7 Limitations
-A main limitation of the scope of our analysis is that due to the significant compute resources required to train LLMs of the size studied in this paper, our exploration is restricted to pretrained LLMs and those fine-tuned on them.
-Several of these models are trained independently of each other, without carefully controlling for different experimental parameters.
-Consequently, direct comparisons between models become challenging, and so the cause of varying levels of regularity between LLMs remains unclear.
-Nonetheless, this makes the consistent emergence of these properties interesting for further study.
-8 Conclusion
-A detailed description of the complexities of transformer blocks in the context of language modelling is an active area of research.
-Our primary goal was to enhance the understanding of the mechanics underlying transformer architectures.
-We provided a quantitative description of embedding trajectories within LLMs.
-In a large collection of models, independently trained by various organizations, we discover a remarkable level of regularity among transformer representations characterized by the alignment of residual Jacobians and emergence of linear and exponentially growing trajectories.
-We observe a correlation between the strength of alignment of residual Jacobian singular vectors and LLM benchmark scores, highlighting the significance of (TA2) and it’s implications regarding performance.
-These findings open avenues for future research for a deeper understanding the connections between regularity of hidden representations and model specifications, LLM architecture, and generalization.
-Appendix A Appendix and Supplementary material
-A.1 Suite of Large Language Models and Prompt Data
-We evaluate 38 total LLMs (24 base LLMs and 14 fine-tuned) on 6 datasets from the HuggingFace Open LLM leaderboard [Beeching et al., 2023].
-Base models.
-Falcon (40B, 7B) [Almazrouei et al., 2023], Llama-3 (70B, 8B) [AI@Meta, 2024], Llama-2 (70B, 13B, 7B) [Touvron et al., 2023], MPT (30B, 7B) [Team, 2023a, b], Mistral v0.1 [Jiang et al., 2023], Gemma (7B, 2B) [Team et al., 2024], Gemma 1.1 (7B, 2B), NeoX (20B) [Black et al., 2022], Neo (2.7B) [Black et al., 2021, Gao et al., 2020], Pythia (6.9B).[Biderman et al., 2023], and GPT-2 (1.5B, 774M, 355M, 117M) [Radford et al., 2019].
-Fine-tuned models.
-CodeLlama (34B, 13B, 7B) [Rozière et al., 2024], CodeLlama Instruct (34B, 13B, 7B) [Rozière et al., 2024], Mistral-v0.1 Instruct (7.3B) [Jiang et al., 2023], Mistral-v0.2 Instruct [Jiang et al., 2023], CodeGemma [Team et al., 2024].
-Datasets.
-ARC [Clark et al., 2018], HellaSwag [Zellers et al., 2019], MMLU [Hendrycks et al., 2021], Truthful QA [Lin et al., 2022], WinoGrande [Sakaguchi et al., 2019].
-Table 1: LLMs featured in the experiments throughout paper.
-Included in the table is the parameter budget of each model, the embedding dimension, the number of training tokens, and the Open LLM leaderboard [Beeching et al., 2023] benchmark score.
-Model Parameters Layers ($L$) Dim. ($d_{\text{model}}$) Training Tokens Score
-Llama-3 70 B 80 8192 15 T
-8 B 32 4096 15 T 62.35
-Llama-2 70 B 80 8192 2 T 66.05
-13 B 40 5120 2 T 55.69
-7 B 32 4096 2 T 50.97
-Codellama 34 B 48 8192 2 T 55.33
-13 B 40 5120 2 T 45.82
-7 B 32 4096 2 T 39.81
-Codellama (IT) 34 B 48 8192 2 T 43.0
-13 B 40 5120 2 T 37.52
-7 B 32 4096 2 T 40.05
-Orca 13 B 40 5120 2 T 58.64
-7 B 32 4096 2 T 54.55
-Falcon 40 B 60 8192 1 T 58.07
-7 B 32 4544 1.5 T 44.17
-Falcon (IT) 40 B 60 8192 1 T 43.26
-7 B 32 4544 1.5 T
-MPT 30 B 48 7168 1 T 66.98
-7 B 32 4096 1 T 56.83
-Phi 2 B 32 2560 1.4 T 61.33
-1.5 B 24 2048 150 B 47.69
-1 B 24 2048 54 B
-Mistral-v0.1 7.3 B 32 4096 60.97
-Mistral-v0.1 (IT) 7.3 B 32 4096 54.96
-Mistral-v0.2 (IT) 7.3 B 32 4096 65.71
-Gemma 7 B 28 3072 6 T 64.29
-2 B 18 2048 6 T 42.75
-Gemma-1.1 7 B 28 3072 6 T 30.0
-2 B 18 2048 6 T 60.09
-Codegemma 7 B 28 3072 6 T 56.73
-2 B 18 2048 6 T 32.19
-Neo 20 B 44 6144 41.69
-2.7 M 32 2560 0.42 36.20
-Pythia 6.9 M 32 4096 0.3 T 39.30
-GPT-2 1.5 B 48 1600 34.12
-774 M 36 1280 32.07
-355 M 24 1024 29.87
-117 M 12 768 28.53
-A.2 Dynamical Motivation
-The equality of top left and right singular vectors (RA2) suggests that the linearizations form a simple dynamical system that acts on representations. Consider a homogenous linear ODE:
+Singular value scaling is more present in untrained transformers (FigureLABEL:fig:svalues_untrained), likely caused by additional layer normalizations in residual blocks (Section 4.1).
+A.4 Dynamical Motivation
+The equality of top left and right singular vectors suggests that the linearizations form a simple dynamical system that acts on representations. Consider a homogenous linear ODE:
 $x^{\prime}=Ax$ (17)
@@
 This agrees with Veit et al.
-[2016] which views ResNets as
+(2016) which views ResNets as
 $x(t)=\left(I+\frac{tA_{1}}{L}\right)\left(I+\frac{tA_{2}}{L}\right)\dots\left(I+\frac{tA_{L}}{L}\right)x(0)$ (19)
-However, Veit et al. [2016] do not make any assumptions about the alignment of the various $A_{i}$ matrices. The Residual Alignment phenomenon suggests ResNets as implementing the simpler system
+However, Veit et al. (2016) do not make any assumptions about the alignment of the various $A_{i}$ matrices. The coupling phenomenon suggests ResNets as implementing the simpler system
 $x(t)=\left(I+\frac{tA}{L}\right)\left(I+\frac{tA}{L}\right)\dots\left(I+\frac{tA}{L}\right)x(0),$ (20)
@@
 Borrowing from the dynamical perspective, we quantify the similarity of hidden trajectories to the evolution of a linear ODE, in order to detect the emergence of a simple linearization to representations.
-A.3 Experimental and Implementation Details
+A.5 Experimental and Implementation Details
 The source code used to produce the results reported in this experiment has been included as supplemental material.
@@
   • $n_{\text{param}}\geq 40B$: 50 prompts per batch.
-Due to the high memory requirement for computing Residual Jacobians, for experiments involving Residual Jacobians, NVIDIA Quadro RTX 6000 was used additionally for $13B>n_{\text{param}}\geq 7B$ and corresponding models were quantized.
-Additionally, due to compute restrictions, alignment of singular vectors of Residual Jacobians was computed on a smaller subset of the 1,200 prompts.
-A.4 Additional Experimental Results
+Due to the high memory requirement for computing Jacobians, for experiments involving the Jacobians, NVIDIA Quadro RTX 6000 was used additionally for $13B>n_{\text{param}}\geq 7B$ and corresponding models were quantized.
+Additionally, due to compute restrictions, alignment of singular vectors of Jacobians was computed on a smaller subset of the 1,200 prompts.
+A.6 Additional Experimental Results
+Figure 15: Hidden Trajectories in LLMs.
+Principal components of the trajectories of the hidden representations through various LLMs (columns, decreasing in model size, see Table 1) in the prompt: What is the capital of France?
+The capital is.
+Top row: all layers.
+Middle Row: layers in shallower transformer blocks (layers specified above plot).
+Bottom Row: layers in deeper transformer blocks (layers specified above plot).
+Trajectories of each input token (last token ‘is’ is plotted in black) are plotted in latent space, visualized with a 2-dimension principal component projection.
+Representations proceed in distinct outward directions, especially in the second half of transformer blocks (lower row) during which the norm of representations increases, with possible abrupt change in the last layer (outer points in upper row).
+A clear direction of movement is visible in each token trajectory.
 Figure 21: Coefficient of variation of layer-wise equidistance.
@@
 The models are sorted by increasing benchmark performance.
+Figure 22: Coefficient of variation of layer-wise expodistance.
+Variation of layer-wise expodistance (Section 3.3) computed over 1,200 prompts from the HuggingFace Open LLM Leaderboard datasets (Section 4.2) on a suite of untrained LLMs (Appendix A.1).
+Figure 23: Various Measurements of Representations.
+All measurements were made on 100 prompts taken from the WikiText 2 datsets.
+(Row 1) Norm of hidden representations as a function of layer depth.
+(Row 2) Line Shape Score (LSS) of the hidden trajectories as a function of layer depth.
+(Row 3) Mean equidistance of contiguous hidden trajectories as a function of depth.
+(Row 4) Entropy of logit vectors as a function of depth.
+Noted in most plots is a line where the behaviour of the measurement drastically changes.
+Figure 24: Linearity Emerges with Training.
+Two plots displaying the evolution of the linearity of the token trajectories through training.
+(Left) The LSS as a function of training checkpoint for the variants of the Pythia Scaling Suite Biderman et al.
+(2023).
+Here, the LSS is measured over each entire prompt.
+(Right) The mean LSS as a function of layer depth measured at various checkpoints throughout the Pythia 12B model.
+Here, the LSS is computed on a window of layers of width $11$, centred at the value given by the x-axis.
+Figure 32: LSS sorted by LLM parameters.

### v3

@@
 In this work, we trace the trajectories of individual tokens as they pass through transformer blocks, and linearize the system along these trajectories through their Jacobian matrices.
-By examining the relationships between these Jacobians, we uncover a transformer block coupling phenomenon in a variety of LLMs, characterized by the coupling of their top singular vectors across tokens and depth.
+By examining the relationships between these Jacobians, we uncover the transformer block coupling phenomenon in a multitude of LLMs, characterized by the coupling of their top singular vectors across tokens and depth.
 Our findings reveal that coupling positively correlates with model performance, and that this relationship is stronger than with other hyperparameters, namely parameter budget, model depth, and embedding dimension.
@@
 1 Introduction
-In recent years, many openly available Large Language Models (LLMs) have been released, achieving state-of-the-art results on task-specific benchmarks.
-The abundance of models, each with differing architecture and training methodology, motivates comparing the underlying mechanisms that drive generalization.
 Transformers (Vaswani et al., 2017) can be represented as discrete, nonlinear, coupled dynamical systems, operating in high dimensions (Greff et al., 2016; Papyan et al., 2017; Haber & Ruthotto, 2017; Ee, 2017; Ebski et al., 2018; Chen et al., 2018; Bai et al., 2019; Rothauge et al., 2019; Gai & Zhang, 2021; Li & Papyan, 2023).
-The term discrete reflects the network’s finite depth; nonlinear refers to the model’s nonlinear components; coupled refers to the interdependent token trajectories enabled by self-attention, and dynamical is due to the residual connections spanning various layers.
-Throughout this work, we investigate whether we can identify structural characteristics that consistently occur in pretrained LLMs that are related to their generalization, as measured on the HuggingFace OpenLLM Leaderboard benchmark.
-To this end, we study and quantify the trajectories of token embeddings and their interactions, through linearizations of LLM layers given by the Jacobians of the associated transformer blocks.
+Viewing the skip connections as enabling a discrete time step, we represent the hidden representations as dynamically evolving through the layers of the network.
+The term nonlinear refers to the nonlinear transformations introduced by activation functions, and coupled refers to the interdependent token trajectories that interact through the MLP and self-attention blocks.
+In our work, we investigate whether there are identifiable structural characteristics across 38+ pretrained LLMs, measure their emergence with training, and analyze their relationship with generalization performance.
+During inference, as token embeddings pass through the network, we linearize the effect of transformer blocks on the token embeddings throughout the depth of the LLM.
+To this end, we compute the Jacobians of distinct connections between layers or tokens, derive their singular value decompositions (SVDs), and compare the resulting singular vectors.
+This approach measures the degree of coupling between singular vectors to capture the operational similarity of blocks as they act on tokens.
 This perspective raises several questions:
 Q1.
-What regularity properties do these trajectories exhibit individually, and what are their relations with one another?
+What regularity properties do these trajectories exhibit, and what are their relations with one another?
 More concretely, what is the relation between the Jacobians across different tokens and transformer layers?
@@
 Q3. Are any of these properties related to the generalization capabilities of LLMs?
-Figure 1: Relationship between Coupling and Benchmark Score.
-The plot illustrates the relationship between average coupling (negative mis-coupling) and benchmark score across various base LLMs (that are not fine-tuned).
-The data shows a noticeable trend: increased alignment correlates with improved benchmark scores.
-The regression fit produces an $R^{2}$ value of $0.7214$.
+Figure 1: Transformer Coupling Measurements.
+(a) The plot illustrates the correlation between average coupling (negative mis-coupling) and benchmark scores across base LLMs (not fine-tuned), showing that higher alignment corresponds to improved performance, with a regression fit yielding an $R^{2}$ value of 0.75 with a significant p-value of $1.56\times 10^{-6}$.
+(b) The mean normalized coupling (Section 3.1) is plotted as a function of training checkpoints for Pythia 12B and 6.9B (Biderman et al., 2023), measured at steps ${128,256,512,1k,2k,\ldots,128k,143k}$.
+(c-e) Adjacency plots illustrate the mean coupling scores between pairs of layers.
+Each node represents a layer, and edge weight and opacity indicate the strength of depth-wise normalized coupling.
+Visualizations are provided for checkpoints 1, 4k, and 143k of Pythia 12B.
 1.1 Contributions
@@
 Coupling.
-The singular vectors of the Jacobians of transformer blocks couple across depth (Figures LABEL:fig:paris_all_8to16, LABEL:fig:paris_all_8to16_appendix) and tokens (Figures LABEL:fig:coupling_inoutsame, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout, LABEL:fig:coupling_inoutsame_appendix, LABEL:fig:coupling_fixed_input_appendix, LABEL:fig:coupling_fixed_output_appendix).
-Further, coupling across Jacobians emerges with training (Figures 2, LABEL:fig:paris_all_8to16, LABEL:fig:coupling_inoutsame, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout).
+The singular vectors of the Jacobians of the transformer blocks are coupled across depth (Figures 3, LABEL:fig:paris_all_8to16_appendix) and tokens (Figures 4, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout, LABEL:fig:coupling_inoutsame_appendix, LABEL:fig:coupling_fixed_input_appendix, LABEL:fig:coupling_fixed_output_appendix) in several open source LLMs (Table (1)).
+Furthermore, coupling across Jacobians emerges with training (Figures 1b, 3, 4, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout), and the coupling strength becomes more pronounced between adjacent layers with training, indicating a layer-wise locality in the interactions (Figures 1(c-e)).
 2.
 Generalization.
-The strength of coupling is correlated with benchmark performance on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) (Figures 1, LABEL:fig:alignment_vary_k).
+The strength of coupling is correlated with benchmark performance on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) (Figures 1a, LABEL:fig:alignment_vary_k).
 Additionally, coupling is more strongly correlated with generalization than parameter budget, model depth, and token embedding dimension (Figure LABEL:fig:correlation_hyperparams).
@@
 Regularity.
-Linearity in hidden trajectories emerges with training (Figures LABEL:figs:lss-training, LABEL:figs:pca, 13, 15, 23, 24).
-Exponential growth occurs in contiguous token representations as a function of depth (Figures 14, 22, 23, LABEL:figs:expo-training).
-We provide a new perspective on token embedding interactions within LLMs by examining transformers layers through their Jacobians.
+Linearity in hidden trajectories emerges with training (Figures LABEL:figs:lss-training, LABEL:figs:pca, LABEL:fig:lss-all, 17, 24, 25), aligning with behaviour previously observed in ResNets Li & Papyan (2023).
+Exponential growth occurs in contiguous token representations as a function of depth (Figures LABEL:fig:expo-all, 23, 24, LABEL:figs:expo-training), starkly contrasting linear growth previously observed.
+We provide a new perspective on token embedding interactions within LLMs by examining layers of a transformer through their Jacobian matrices.
 Our results display the effect of training on transformer blocks, and suggest potential approaches for promoting generalization in LLMs.
-Figure 2: Emergence of Singular Vector Coupling with Training.
-Normalized coupling (Section 3.1) evaluated for $J_{i,j}^{d_{1}},J_{i,j}^{d_{2}}$, with $|d_{1}-d_{2}|\leq 5$.
-Measurements are taken for checkpoints of Pythia 12B, Pythia 6.9B, Pythia 1B (Biderman et al., 2023) at training steps $\{1,16,128,13\text{k},23\text{k},33\text{k},\ldots,143\text{k}\}$.
-Coupling is low at early training iterations, and increases with training, with Pythia 12B and 6.9B displaying greater coupling than Pythia 1B in final training steps.
 2 Background on Large Language Models
 We describe LLMs as a deep composition of functions that iteratively transform token embeddings.
-In the input layer, $l=0$, textual prompts undergo tokenization and are combined with positional encodings to create an initial high-dimensional embedding, denoted by $x_{i}^{0}\in\mathbb{R}^{d_{\text{model}}}$ for the $i^{\text{th}}$ token.
+In the input layer, $l=0$, the text prompts are tokenized and combined with the positional encodings to create an initial high-dimensional embedding, denoted by $x_{i}^{0}\in\mathbb{R}^{d_{\text{model}}}$ for the $i^{\text{th}}$ token.
 When these embeddings are stacked, they form a matrix:
@@
 $F_{\text{block}}^{L}(X^{L-1})$ $=\text{LN}(X^{L-1}+h^{L}(X^{L-1})+\text{FFN}(g^{L}(X^{L-1}))).$ (7)
-The output $X^{L}$ from the final block $F^{L}$ is passed into a bias-free linear layer $M\in\mathbb{R}^{d_{\text{vocab}}\times d_{\text{model}}}$, with $d_{\text{vocab}}$ denoting the size of the token vocabulary and $d_{\text{model}}$ is the dimension of the token embeddings.
+The output $X^{L}$ of the final block $F^{L}$ is passed into a bias-free linear layer $M\in\mathbb{R}^{d_{\text{vocab}}\times d_{\text{model}}}$, with $d_{\text{vocab}}$ denoting the size of the token vocabulary and $d_{\text{model}}$ is the dimension of the token embeddings.
 This layer $M$ computes final-layer logits for each token embedding, $\ell_{i}=Mx_{i}^{L}$.
 The prediction for the next token is then determined by selecting the maximal logit value: $\arg\max_{v\in\text{tokens}}\ell_{v,n}.$
+Figure 2: Transformer Block Coupling.
+A visualization of the various types of transformer block coupling with brief instructions on computing both the Jacobians $J$ and coupling matrices $A$ (Section 3.1).
+The coupling measurement quantifies the alignment and agreement between the interactions of embeddings connections within the network.
+The colored subscripts in the sample matrices $A$ indicate the specific connections being compared.
 3 Methods
 3.1 Coupling of Singular Vectors of Jacobians
-Jacobians. Coupling is investigated through analyzing the linearizations of transformer blocks given by their Jacobians
+Jacobians. Coupling is investigated through analyzing the linearizations of transformer blocks which is given by their Jacobian matrices
 $J_{t_{1}t_{2}}^{l}=\frac{\partial}{\partial x_{t_{1}}^{l-1}}\left(f^{l}(X^{l-1})\right)_{t_{2}},$ (8)
 defined for each layer $l\in\{1,\ldots,L\}$, and pair of tokens $t_{1},t_{2}\in\{1,\ldots,n\}$.
-Note that this is the Jacobian matrix for each transformer block without the contribution from the skip connection from the input of the block, similar to the quantity measured by Li & Papyan (2023), but extended for cross-token interactions.
-Also, due to the causal structure of the representations, $J_{t_{1}t_{2}}^{l}=0$ if $t_{1}>t_{2}$, so we restrict our attention to the case $t_{1}\leq t_{2}$.
-Singular value decomposition.
-The SVD of the Jacobians $J_{t_{1}t_{2}}^{l}$ are computed, i.e., $J_{t_{1}t_{2}}^{l}=U_{l}S_{l}V_{l}^{\top}$ (with superscripts $t_{1},t_{2}$, indicating the tokens, omitted for clarity), where $U_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ and $V_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ are matrices of left and right singular vectors respectively, and $S_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ is the singular value matrix.
-To measure coupling between Jacobians $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}}$, define $A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime},K}\coloneq U_{l^{\prime},K}^{\top}J_{t_{1}t_{2}}^{l}V_{l^{\prime},K}$ for $l,l^{\prime}\in\{1,\ldots,L\}$ and $t_{1},t_{2},t_{1}^{\prime},t_{2}^{\prime}\in\{1,\ldots,n\}$, where $U_{l^{\prime},K}$ and $V_{l^{\prime},K}$ are the sub-matrices with columns that are the top-$K$ left and right singular vectors of $J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}}$, respectively.
-Figure 3: Transformer Block Coupling.
-Displayed is a visualization of the various types of transformer block coupling, along with brief instructions on computing both the Jacobians $J$ and coupling matrices $A$.
-Depth-wise Coupling.
-To analyze coupling across transformer blocks, we fix $t$ (usually as the last token), and check alignment between $J_{tt}^{l}$ and $J_{tt}^{l^{\prime}}$ through $A_{t}^{ll^{\prime},K}$ (we write the single subscript $t$ for clarity) across layers $l,l^{\prime}\in\{1,\ldots,L\}$.
-Token-wise Coupling.
-To study the coupling across tokens, we start by fixing two layers $l,l^{\prime}\in\{1,\ldots,L\}$, and analyze coupling across different cases of the $t_{1},t_{2},t_{1}^{\prime},t_{2}^{\prime}$.
-We start with the case where the input and output tokens are the same and compare $J_{tt}^{l}$ and $J_{t^{\prime}t^{\prime}}^{l^{\prime}}$ across $t,t^{\prime}\in\{1,\ldots,n\}$, which is the coupling across tokens for a token’s effect on its own trajectory.
-We also consider the context tokens’ impact on a trajectory by including coupling across cases with $t_{1}<t_{2}$.
-More specifically, we measure coupling between $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}t_{2}^{\prime}}^{l^{\prime}}$ across $t_{2},t_{2}^{\prime}\geq t_{1}$ (fixing the input token to be the same) and also between $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}^{\prime}t_{2}}^{l^{\prime}}$ across $t_{1},t_{1}^{\prime}\leq t_{2}$ (fixing the output token to be the same).
-Coupling measurement.
-To quantify the coupling of singular vectors of the Jacobians, we measure how far the $A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime},K}$ are from being diagonal.
-More precisely, we quantify the mis-coupling of $A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime},K}$ using $m:\mathbb{R}^{K\times K}\to\mathbb{R}$ defined by
+Note that this is the Jacobian matrix for each transformer block without the contribution of the skip connection from the input of the block, similar to the quantity measured by Li & Papyan (2023) which strictly analyzes the case where $t_{1}=t_{2}$.
+Due to the causal structure of the representations, $J_{t_{1}t_{2}}^{l}=0$ whenever $t_{1}>t_{2}$. Hence, we restrict our attention to the case where $t_{1}\leq t_{2}$.
+Singular value decomposition (SVD). We compute the SVD of the Jacobians:
+$J_{t_{1}t_{2}}^{l}=U_{l}S_{l}V_{l}^{\top}$
+where $U_{l},V_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ are the matrices of left and right singular vectors respectively, and $S_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ contains the singular values.
+[^1] Note that the superscripts $t_{1},t_{2}$, indicating the tokens, are omitted for clarity in the expression for the SVD.
+Coupling Measurement. To measure coupling between two Jacobians
+$J_{t_{1}t_{2}}^{l}=U_{l}S_{l}V_{l}^{\top}$ and $J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}}=U_{l^{\prime}}S_{l^{\prime}}V_{l^{\prime}}^{\top},$ (9)
+we define the coupling matrix
+$A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime}}$ :-$\coloneq U_{l^{\prime}}^{\top}J_{t_{1}t_{2}}^{l}V_{l^{\prime}}:$ (10)
+$=U_{l^{\prime}}^{\top}U_{l}S_{l}V_{l}^{\top}V_{l^{\prime}}$ (11)
+for $l,l^{\prime}\in\{1,\ldots,L\}$ and $t_{1},t_{2},t_{1}^{\prime},t_{2}^{\prime}\in\{1,\ldots,n\}$. If the singular vectors of distinct Jacobians are strongly aligned, then
+$U_{l^{\prime}}^{\top}U_{l}\approx I\approx V_{l}^{\top}V_{l^{\prime}},$ (12)
+implying that the coupling matrix $A$ should be strongly diagonal. Explicitly, we quantify the mis-coupling of $A$ using
 $m(A)=\lVert A-\text{Diag}(A)\rVert_{F},$ (13)
 where $\text{Diag}(A)$ is the matrix $A$ with all non-diagonal entries replaced by zero and $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm.
-For normalized comparison between models across training in Figure 2, we normalize $A$ by their norms
-$\tilde{m}(A)=\frac{\lVert A-\text{Diag}(A)\rVert_{F}}{\lVert A\rVert_{F}}$ (10)
+For normalized comparison between models, we normalize by $A$;
+$\tilde{m}(A)=\frac{\lVert A-\text{Diag}(A)\rVert_{F}}{\lVert A\rVert_{F}}.$ (14)
+Depth-wise Coupling.
+To analyze coupling across transformer blocks, we fix $t$, and measure alignment between $J_{tt}^{l}$ and $J_{tt}^{l^{\prime}}$ through the matrix $A_{t}^{ll^{\prime}}$ across layers $l,l^{\prime}\in\{1,\ldots,L\}$  [^2] In the matrix $A$, we write the single subscript $t$ for clarity.
+Token-wise Coupling We also quantify the coupling across tokens in several ways:
+  • Self-coupling.
+By fixing two layers $l,l^{\prime}\in\{1,\ldots,L\}$, we analyze the case where the input and output tokens are the same.
+Explicity, we compare $J_{tt}^{l}$ and $J_{t^{\prime}t^{\prime}}^{l^{\prime}}$ across $t,t^{\prime}\in\{1,\ldots,n\}$, which represents the coupling across tokens for a token’s effect on its own trajectory.
+  • Context Coupling.
+We consider the context tokens’ impact on a trajectory by measuring coupling between $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}t_{2}^{\prime}}^{l^{\prime}}$ across $t_{2},t_{2}^{\prime}\geq t_{1}$ (fixing the input token to be the same) and also between $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}^{\prime}t_{2}}^{l^{\prime}}$ across $t_{1},t_{1}^{\prime}\leq t_{2}$ (fixing the output token to be the same).
 3.2 Linearity of Trajectories
@@
 3.3 Layer-Wise Exponential Growth
-We measure the presence of exponential spacing (expodistance) of the hidden trajectories. Assuming exponential growth of the embedding norms as they flow through the hidden layers, we can estimate,
-$\|x_{i}^{l}\|$ $\approx e^{\alpha l}\|x_{i}^{0}\|=e^{\alpha}\|x_{i}^{l-1}\|$ (13)
-for some fixed $\alpha\in\mathbb{R}$ over all layers $l=1,\ldots L$.
-We quantify the validity of the representation in Equation 13 by measuring the coefficient of variation of $\alpha_{i}^{l}$, given by
+We measure the presence of exponential spacing (expodistance) of the hidden trajectories.
+Assuming exponential growth of the embedding norms as they flow through the hidden layers, we estimate $\|x_{i}^{l}\|\approx e^{\alpha l}\|x_{i}^{0}\|=e^{\alpha}\|x_{i}^{l-1}\|$ for some fixed $\alpha\in\mathbb{R}$ over all layers $l=1,\ldots L$.
+We quantify the validity of this representation by measuring the coefficient of variation of $\alpha_{i}^{l}$, given by
 $\alpha_{i}^{l}\approx\ln\left(\frac{\|x_{i}^{l}\|}{\|x_{i}^{l-1}\|}\right),$ (17)
@@
 $\text{ED}_{i}=\frac{\text{Var}_{l}\alpha_{i}^{l}}{(\text{Avg}_{l}\alpha_{i}^{l})^{2}}.$ (18)
-This measurement is motivated by the parametrization discussed in Appendix A.4, as well as empirical evidence in Figure LABEL:figs:norm-expo-traininga, and serves as a method to test the validity of the linearization presented in Equation 17.
+This measurement is motivated by the discussion in Section 6.1 the parametrization discussed in Appendix A.5, as well as empirical evidence in Figure LABEL:figs:norm-expo-traininga, and serves as a method to test the validity of the linearization presented in Equation 19.
+3.4 Vision Transformer Training
+For further investigation of coupling in transformers, we train a series of Vision Transformers (ViTs) following DEiT training (Touvron et al., 2021).
+We train 64 ViTs on CIFAR10 (Krizhevsky, 2009) with varied weight decay and stochastic depth rate for a fixed architecture of embedding size 192, depth 12, and 3 attention heads.
+Please see Appendix A.7 for further details.
+Figure 3: Transformer Block Coupling across Depth.
+The figure shows Jacobian coupling across transformer blocks 9 to 16, using the prompt ”What is the capital of France?
+The capital is” to trace the final token’s trajectory.
+In trained models (bottom row), the diagonal pattern with minimal off-diagonal values indicates alignment of Jacobians, where top singular vectors of $J^{l^{\prime}}$ diagonalize $J^{l}$.
+Untrained models (top row) lack this alignment.
+Further details are in the Appendix.
+A.8 (Figure LABEL:fig:paris_all_8to16_appendix).
+Best viewed in color.
 4 Evaluation
@@
 5.1 Coupling of Jacobians Across Depth
-In trained LLMs, we observe coupling of the top singular vectors of the Jacobians across depth (Figure LABEL:fig:paris_all_8to16 bottom row), evident in the low non-diagonal values with a visible diagonal present in the matrix subplots.
+In trained LLMs, we observe coupling of the top singular vectors of the Jacobians across depth (Figure 3 bottom row), evident in the low non-diagonal values with a visible diagonal present in the matrix subplots.
 This is consistently observed across various LLMs considered.
-On the other hand, in untrained models (Figure LABEL:fig:paris_all_8to16 top row), there is no coupling of Jacobians across different depths.
+On the other hand, in untrained models (Figure 3 top row), there is no coupling of Jacobians across different depths.
 There is coupling along the diagonal, however, because each Jacobian is trivially diagonalized by its own singular vectors.
@@
 5.2 Coupling of Jacobians Across Tokens
-We also analyze the coupling of singular vectors of the Jacobians across tokens.
-First, we observe coupling between Jacobians where the input and output tokens are the same, that is, between $J_{l}^{tt}$ and $J_{l^{\prime}}^{t^{\prime}t^{\prime}}$ across $t,t^{\prime}\in\{1,\ldots,n\}$ (Figure LABEL:fig:coupling_inoutsame).
-This suggests that a token’s interaction on its own trajectory is coupled with that of other tokens.
-We also consider the context tokens’ interaction on a trajectory when fixing the input token $t_{1}$ (Figure LABEL:fig:coupling_fixedin), through $J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}t_{2}^{\prime}}$ across $t_{2},t_{2}^{\prime}\geq t_{1}$, and when fixing the output token $t_{2}$ (Figure LABEL:fig:coupling_fixedout), with $J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}^{\prime}t_{2}}$ across $t_{1},t_{1}^{\prime}\leq t_{2}$.
-Although generally present, the coupling in these two cases are less consistent, with variation in the strength of coupling between different token pairs.
-Again in untrained models (Figure LABEL:fig:coupling_inoutsame top row), there is a lack coupling of Jacobians across tokens.
+We analyze the coupling of singular vectors of Jacobians across tokens.
+For input and output tokens that are the same ($J_{l}^{tt}$ and $J_{l^{\prime}}^{t^{\prime}t^{\prime}}$, Figure 4), we observe strong coupling, indicating that a token’s interactions along its trajectory are coupled with others.
+For context tokens, coupling is examined by fixing the input token ($J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}t_{2}^{\prime}}$, Figure LABEL:fig:coupling_fixedin) or the output token ($J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}^{\prime}t_{2}}$, Figure LABEL:fig:coupling_fixedout).
+While coupling exists, it is less consistent across pairs.
+Untrained models (Figure 4 top row) show no such coupling.
+Figure 4: Transformer Block Coupling across Tokens (Same input and output tokens).
+The figure shows Jacobian coupling for the same input and output token across tokens, visualized using the absolute values of $A_{ll^{\prime}}^{ttt^{\prime}t^{\prime}}$ (with fixed layers $l,l^{\prime}$).
+In trained models (bottom row), the strong diagonal and small off-diagonal values indicate coupling, while no such coupling is present at initialization (top row).
+Additional details are in Appendix A.8 (Figure LABEL:fig:coupling_inoutsame_appendix).
 5.3 Emergence of Coupling with Training
-Coupling emerges through training for the evaluated LLMs, including coupling across depth (Figure LABEL:fig:paris_all_8to16 ) and across tokens (Figures LABEL:fig:coupling_inoutsame, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout).
-Further, we evaluate layer-wise coupling at intermediate training checkpoints of Pythia 12B, 6.9B, and 1B (Biderman et al., 2023) (Figure 2), and observe that coupling is generally low at initialization and increases significantly during early training steps.
+Coupling emerges through training for the evaluated LLMs, including coupling across depth (Figure 3 ) and across tokens (Figures 4, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout).
+Further, we evaluate layer-wise coupling at intermediate training checkpoints of Pythia 6.9B and 12B (Biderman et al., 2023) (Figure 1b), and observe that coupling is generally low at initialization and increases persistently throughout training.
+Moreover, there is a clear sense of locality in the strength of coupling which is visually displayed in Figure (1c-e)
+The gradual growth of coupling observed in Figure (1b) parallels the logarithmic increase in accuracy during training for Pythia 12B and Pythia 6.9B (Biderman et al., 2023).
+This highlights the relationship between coupling and performance, since both properties emerge at similar training iterations and rates.
 5.4 Correlation with Generalization
@@
 This observation suggests a compelling relationship between stronger coupling of singular vectors of Jacobians $J_{l}^{t_{1}t_{2}}$ and improved generalization.
+The results for ViTs demonstrate that stochastic depth encourages coupling during training (Figure LABEL:figure:coupling_vitb) and that coupling correlates with accuracy when fixing SD rate (Figure LABEL:figure:coupling_vita).
+This finding suggests that coupling may provide new insight into stochastic depth’s underlying mechanism, and that developing training methods to amplify coupling across transformer blocks could provide additional regularization and improve performance.
+We hypothesize that simple trajectories may lead to better generalization.
+This agrees with many generalization bounds in machine learning (Arora et al., 2018), which suggest that models with lower complexity tend towards better generalization.
+Additionally, prior works (Novak et al., 2018) demonstrate that the Frobenius norm of input-output Jacobians is related to generalization, providing evidence that coupling — a structural property derived from Jacobians — may also correlate with generalization.
 5.5 Regularity in Hidden Trajectories
 We identify a considerable degree of linearity in the hidden trajectories of all featured LLMs.
-The LSS of the trajectories has an average value $4.25$ in trained models, while taking average values of $6.54$ at initialization, and is supported by the low variation across benchmark prompts (Figure 13).
+The LSS of the trajectories has an average value $4.25$ in trained models, while taking average values of $6.54$ at initialization, and is supported by the low variation across benchmark prompts (Figure LABEL:figs:lss-expo).
 Linearity increases with training at varying depths of Pythia 12B (Figure LABEL:figs:lss-training).
-Linear and expansive behavior of the representations is demonstrated in Llama-3 70B, Mistral 30B, and NeoX 20B through low dimensional projections of embedding trajectories (Figure LABEL:figs:pca).
-Figure 13: Linearity of Trajectories.
-The figure depicts the line-shape score (LSS) of embedding trajectories, as discussed in Section 5.5), computed on 1,200 prompts of the HuggingFace Open LLM Leaderboard (Section 4.2) for a variety of trained (black) and randomly initialized (red) LLMs (Appendix A.1).
-Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model.
-Models are sorted by number of parameters.
-5.6 Layer-Wise Exponential Growth
-Across all LLMs considered, the hidden trajectories exhibit exponential growth that emerges with training (Figure 14).
-The majority of the models exhibit a low coefficient of variation across prompts, showing the robustness of this property across a variety of tasks.
-The expodistance measurements show little dependence on model parameters, and take on consistently low values across all featured LLMs (Figure 14).
+Linear and expansive behavior of the representations is demonstrated in Llama-3 70B, MPT 30B, and NeoX 20B through low dimensional projections of embedding trajectories (Figure LABEL:figs:pca).
+Among the LLMs considered, the vast majority of hidden trajectories exhibit exponential growth that emerges with training (Figure LABEL:fig:expo-all b).
+Many models exhibit a low coefficient of variation across prompts, showing the robustness of this property across a variety of tasks.
 In contrast, measurements at initialization show equally (rather than exponentially) distanced trajectories, as reflected by the low coefficients of variation in the norms of their layer-wise differences (Figure 23).
-Under certain assumptions, exponential spacing is justified in Section 6.2.
-Figure 14: Coefficient of variation of layer-wise exponent.
-The figure presents the coefficient of variation of the layer-wise exponent (Section 3.3) computed on 1,200 prompts of the HuggingFace Open LLM Leaderboard (Section 4.2) for trained LLMs (Appendix A.1).
-Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model.
-LLMs are sorted by increasing benchmark performance, while the vertical axis has a base 10 logarithmic scale.
+Under certain assumptions, exponential spacing is motivated in Section 6.1.
 6 Discussion
-6.1 Correlation with Generalization
-As discussed in Section 5.4, stronger coupling across Jacobian singular vectors is related to improved generalization (Figure 1).
-This relationship may assist in evaluating the training of transformers.
-In particular, this finding suggests that the development of training methods to amplify coupling across transformer blocks may lead to favourable model performance.
-6.2 Emergence of Exponential Spacing with Coupling
-The emergence of exponential spacing in many LLMs can be viewed as a result of coupling under some assumptions on the homogeneity of the singular values across depth.
-Strong coupling suggests that the linearizations of hidden trajectories satisfy a linear difference equation (see Appendix A.4), given by
-$x^{l+1}-x^{l}=Ax^{l}$
-Considering the singular value decomposition of $A$ in Equation 18 as $A=U\Sigma V^{T}$, and assuming that there is coupling ($U=V$), the difference equation can be diagonalized in the same basis across depth.
-Further assuming that the top singular values of the Jacobian blocks display low variation across depth, (Figure LABEL:fig:svalues_trained), we expect
-$||x^{l}||\approx e^{\sigma l}\,||x^{0}||,$ (16)
-where $\sigma$ is the top singular value.
-Since we assume $\sigma$ is independent of depth, the above norm is expected to be exponential in $l$, and offers an explanation to the measurements in Figure 14.
+6.1 Emergence of Regularity with Coupling
+Under certain assumptions, the emergence of increased linearity and exponential spacing in many LLMs can be analyzed as a result of coupling.
+Considering input embeddings $x_{1}^{0},\ldots,x_{n}^{0}$ and the linearization of the last token embedding $x_{n}^{l}$ given by $J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})$:
+$x_{n}^{l}=(I+J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0}))x_{n}^{l-1}$
+We simplify notation and write $x_{n}^{l}=x^{l},\,J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})=J^{l}$. The representations follow the linearized equation:
+$x^{l+1}=x^{l}+J_{l}x^{l}=(I+J_{l})x^{l}.$
+Expanding across layers, the entire system can be approximated by the product
+$x^{L}=(I+J_{L})(I+J_{L-1})\cdots(I+J_{1})x^{0}.$
+Assuming that $J_{l}\approx USU^{T}$, this equation predicts that the norm of $x_{l}$ would exhibit exponential growth layer by layer. Expanding $U$ and $S$,
+$x^{l}=\sum_{j=1}^{d_{\text{model}}}u_{j}(1+s_{j})^{l}u_{j}^{\top}x^{0}.$
+where $u_{j}$ and $s_{j}$ represent the eigenvectors and eigenvalues of the Jacobian, respectively.
+In general, trajectories are not expected to be perfectly linear unless $x^{0}$ aligns with an eigenvector of $J$.
+However, in our experiments, we observe a notable tendency towards linearity, suggesting that the representations align progressively during training with the eigenvectors of the coupled Jacobians.
+6.2 Significance of Coupling
+The coupling phenomenon provides insight into the internal operations of transformer.
+We hypothesize that during training, the LLM learns to represent embeddings in specific low-dimensional subspaces (Eldar & Mishali, 2009).
+Given an input, the first layer converts the input into embeddings within one of these learned subspaces.
+Each subsequent transformer layer modifies these embeddings, potentially moving them to different subspaces.
+Strong coupling between consecutive layers suggests that the LLM tends towards representations in the same or similar subspaces across many layers.
+Weak coupling suggests that the subspaces may change between layers, though usually gradually, and that adjacent layers still operate in relatively similar subspaces.
+Previous works have shown (Lad et al., 2024; Gromov et al., 2024) that the early and late layers of language models behave differently, which may be understood through coupling; tokens remain in similar spans, then transition to a different subspace, continuing within a new span that is consistent in the remainder of the transformer.
+The emergence of coupling with training steps (Figure 1) may provide insight into the dynamics. Under full coupling and a difference equation approximation, the representations evolve as
+$x^{l}=\sum_{j=1}^{d_{\text{model}}}u_{j}(1+s_{j})^{l}u_{j}^{\top}x^{0}$
+where $u_{i}$ and $\lambda_{i}$ denote the eigenvectors and eigenvalues of the Jacobian, respectively.
+In this case, the gradients of the loss $L$ with respect to prediction $y$, $x^{L}$ are represented by
+$\frac{\partial\text{L}}{\partial x^{0}}(x^{L},y)\approx\sum_{j=1}^{d_{\text{model}}}u_{j}(1+s_{j})^{L}u_{j}^{\top}(y-x^{L}),$
+Due to the coupling, the dynamics exhibit either exponential growth or decay in different subspaces, depending on the sign of $s_{j}$, which is known to cause challenges for optimization as in past works on dynamical isometry (Pennington et al., 2017).
+We infer that increasing coupling during training makes optimization progressively more difficult.
+Conversely, as training progresses, it becomes harder to achieve stronger coupling, and is consistent with the logarithmic trend in Figure 1.
 7 Related Work
@@
 We show coupling in $J_{l}^{t_{1}t_{2}}$ across various tokens, which is specific to tokenization in transformers, in addition to demonstrating coupling of $J_{l}^{t_{1}t_{2}}$ across $l$, which was also identified analogously in ResNets (Li & Papyan, 2023).
-Our measurements quantify the emergence with training of linearity and spacing, including at different depths, which was previously observed mostly qualitatively.
 Further comparison is included in Appendix A.3.
-Neural ODE.
-Introduced by (Chen et al., 2018), Neural ODE view ResNets as a discretized dynamical process, with past work (Sander et al., 2022) showing the convergence of Residual Networks to a system of linear ODE.
-(Zhong et al., 2022; Li et al., 2021) studied neural ODE specific perspective to transformers and other LLM architectures.
-In our work, the presence of residual connections and emergence of Jacobian Alignment in transformers suggests that a discretization of a simple ODE emerges in LLMs.
-Information Processing in LLMs.
-Previously, Katz & Belinkov (2023); Bietti et al.
-(2023) have described memory and semantic flow in intermediate states of transformers while van Aken et al.
-(2019) have examined BERT hidden trajectories.
-Investigation of LLMs for reasoning tasks and interpretability remains an important research focus (Huang & Chang, 2023).
+Neural Ordinary Differential Equations.
+Neural ODEs (Chen et al., 2018) view ResNets as a discretized dynamical process, with past work (Sander et al., 2022) showing the convergence of Residual Networks to a system of linear ODE, with some extensions to transformers (Zhong et al., 2022; Li et al., 2021) The emergence of coupling in transformers suggests that a discretization of a simple itertaive process emerges in LLMs.
 In-context learning.
-Previous works have observed that LLMs are able to perform in context learning (von Oswald et al., 2023; Bai et al., 2024; Ahn et al., 2023; Akyürek et al., 2023; Xie et al., 2021; Hahn & Goyal, 2023; Xing et al., 2024): the model is capable of performing a particular machine learning task through a few examples provided in a single prompt or context.
-Their results consider a single layer of trained self-attention as implementing a single step of a gradient-descent-like learning rules across transformers depth, minimizing the MSE of a linear model
-$x_{\text{final}}=\min_{x}||Ax-b||^{2}$
-A gradient-descent learning rule would cause depth-wise updates to representations of the form
-$x_{t+1}=x_{t}+\epsilon A^{T}(Ax_{t}-b)=(I+\epsilon A^{T}A)x_{t}-{\epsilon}A^{T}b$
-We should expect similarity in the matrices $I+{\epsilon}A^{T}A$ across depth, which is suggested by coupling.
+LLMs can perform tasks through examples provided in a single prompt, demonstrating in-context learning (von Oswald et al., 2023; Bai et al., 2024; Ahn et al., 2023; Akyürek et al., 2023; Xie et al., 2021; Hahn & Goyal, 2023; Xing et al., 2024).
+Studies suggest trained self-attention layers implement gradient-descent-like updates across depth to minimize the MSE of a linear model:
+$x_{\text{final}}=\min_{x}||Ax-b||^{2}.$
+These updates take the form:
+$x_{t+1}=(I+\epsilon A^{T}A)x_{t}-\epsilon A^{T}b.$
+Coupling across depth suggests similarity in the matrices $I+\epsilon A^{T}A$.
 Hidden Representation Dynamics.
-Previous studies (Gai & Zhang, 2021; Haber & Ruthotto, 2017; Ee, 2017), have focused on interpreting deep neural networks from a dynamical systems perspective.
-Throughout training, numerical experiments have shown that both plain networks and ResNets are aptly approximated by a geodesic curve in Wasserstein space (Gai & Zhang, 2021).
-Other interpretations involve partitioning activation space into basins of attraction, with fixed points representing discrete thoughts (Nam et al., 2023).
+Prior research interprets deep neural networks through dynamical systems, revealing that training trajectories align with geodesic curves (Gai & Zhang, 2021) and partition activation space into basins of attraction (Nam et al., 2023).
 For further related works, see (Geshkovski et al., 2023b; a; Tarzanagh et al., 2023; Valeriani et al., 2023).
 Structure in Hidden Representations.
-Emergence of regularity in last-layer representations during training was noticed in Neural Collapse (Papyan et al., 2020).
-Subsequent works examined structure of hidden representations and presented theoretical results describing the emergence of Neural Collapse (Wang et al., 2024a; Parker et al., 2023; Zangrando et al., 2024; Garrod & Keating, 2024; Wang et al., 2024b; Hoyt & Owen, 2021; Arous et al., 2023; Zarka et al., 2021; Ben-Shaul & Dekel, 2022; Papyan, 2020; Súkeník et al., 2023).
-Within LLMs, recent works have identified uniform structure in token means (Wu & Papyan, 2024), and low dimensional structure across hidden trajectories (Sarfati et al., 2024).
-Our work studies the representations of all layers in LLMs in a local sense through the interactions of tokens through their Jacobians.
+Neural Collapse (Papyan et al., 2020) highlights emergent regularity in last-layer representations, with subsequent studies exploring hidden-layer structures and their theoretical underpinnings (Wang et al., 2024a; Parker et al., 2023; Zangrando et al., 2024; Garrod & Keating, 2024; Wang et al., 2024b; Hoyt & Owen, 2021; Arous et al., 2023; Zarka et al., 2021; Ben-Shaul & Dekel, 2022; Papyan, 2020; Súkeník et al., 2023).
+In LLMs, recent works identify uniform token structures (Wu & Papyan, 2024) and low-dimensional hidden trajectories (Sarfati et al., 2024).
+Our work examines local token interactions through Jacobian dynamics across all LLM layers.
 8 Limitations
-A limitation of the scope of our analysis is that due to the significant compute resources required to train LLMs of the size studied in this paper, our exploration is restricted to pretrained LLMs and those fine-tuned on them.
-Several of these models are trained independently of each other, without carefully controlling for different experimental parameters.
-Consequently, direct comparisons between models become challenging, and so determining the cause of varying levels of regularity between LLMs becomes difficult.
-Nonetheless, this makes the consistent emergence of these properties interesting for further study.
+Our analysis is limited to pretrained LLMs and their fine-tuned variants due to the high computational cost of training large models.
+Variations in experimental setups across independently trained models hinder direct comparisons, making it challenging to pinpoint causes of differing regularity.
+However, the consistent emergence of these properties warrants further study.
 9 Conclusion
@@
 Our research builds on the understanding of transformer architectures by revealing the coupling, across depth and token, of singular vectors in the Jacobians of transformer blocks for multiple LLMs trained by various organizations.
-We additionally observe and quantify the emergence of linearity and exponential spacing with training in the embedding trajectories of LLMs.
-Notably, we establish a correlation between the strength of this coupling and benchmark performance on the HuggingFace Open LLM Leaderboard, highlighting the significance of transformer block coupling for generalization.
+We establish a correlation between the strength of this coupling and benchmark performance on the HuggingFace Open LLM Leaderboard, highlighting the significance of transformer block coupling for generalization.
 These findings open avenues for future research, encouraging deeper exploration into the connections between regularity of hidden representations, model specifications, and generalization.
@@
 Linearity in hidden trajectories, as observed in ResNets Li & Papyan (2023); Gai & Zhang (2021), also emerges with training in LLMs.
-The mean LSS value among the evaluated LLMs is $4.24$ (Figure 13), greater than LSS measurements observed for ResNets (Gai & Zhang (2021), page 18) which range between $2.0$-$3.0$ (due to varying trajectory length and hidden dimension).
+The mean LSS value among the evaluated LLMs is $4.24$ (Figure LABEL:fig:lss-all), greater than LSS measurements observed for ResNets (Gai & Zhang (2021), page 18) which range between $2.0$-$3.0$ (due to varying trajectory length and hidden dimension).
 In both architectures, training induces improved linearity and regularity (Figure LABEL:figs:pca) in trajectories.
-In contrast to ResNets, trajectories are not equidistant, instead showing exponential growth between layers (Figure 14).
+In contrast to ResNets, trajectories are not equidistant, instead showing exponential growth between layers (Figure LABEL:fig:expo-all).
 We quantify this spacing through a low coefficient of variation, displaying the presence of exponential growth in token trajectories.
@@
 Singular value scaling is more present in untrained transformers (FigureLABEL:fig:svalues_untrained), likely caused by additional layer normalizations in residual blocks (Section 4.1).
-A.4 Dynamical Motivation
-The equality of top left and right singular vectors suggests that the linearizations form a simple dynamical system that acts on representations. Consider a homogenous linear ODE:
-$x^{\prime}=Ax$ (17)
-Its solution is given by
-$x(t)=\exp(tA)x(0)=\lim_{L\to\infty}\left(I+\frac{At}{L}\right)^{L}x(0)$ (18)
-where the product is repeated $L$ times.
-Expanding the brackets shows that $x(t)$ can be thought of as a collection of many paths of various lengths, due to the binomial identity.
-This agrees with Veit et al.
-(2016) which views ResNets as
-$x(t)=\left(I+\frac{tA_{1}}{L}\right)\left(I+\frac{tA_{2}}{L}\right)\dots\left(I+\frac{tA_{L}}{L}\right)x(0)$ (19)
-However, Veit et al. (2016) do not make any assumptions about the alignment of the various $A_{i}$ matrices. The coupling phenomenon suggests ResNets as implementing the simpler system
-$x(t)=\left(I+\frac{tA}{L}\right)\left(I+\frac{tA}{L}\right)\dots\left(I+\frac{tA}{L}\right)x(0),$ (20)
-where all the $A_{i}$ matrices are aligned. One benefit of this interpretation is that, as $L\to\infty$, we can write $x(t)$ in a simple closed form, namely $x(t)=\exp(tA)x(0)$.
-Borrowing from the dynamical perspective, we quantify the similarity of hidden trajectories to the evolution of a linear ODE, in order to detect the emergence of a simple linearization to representations.
-A.5 Experimental and Implementation Details
+A.4 Significance of Coupling
+The transformer block coupling phenomenon offers insight into several prominent practices in LLM research, as summarized in Table 2.
+Table 2: Significance of the Coupling Phenomenon. A table which highlights the implications of transformer block coupling to a variety of effors in machine learning research.
+Current Research Practice Key Idea Our Contribution
+Compressing models by merging blocks (Fu et al., 2022; Kim et al., 2024) Combine adjacent transformer blocks to reduce model size without significant performance loss.
+Demonstrates that merging is effective because blocks become strongly coupled during training.
+Compressing models by pruning blocks (Elkerdawy et al., 2020; Kim et al., 2024; Dror et al., 2021; Fang et al., 2023) Remove certain transformer blocks while preserving functionality.
+Explains that pruning works because the coupling ensures redundancy across blocks.
+Compressing models by projecting weight matrices (Ashkboos et al., 2024) Reduce dimensionality by projecting weights into smaller subspaces.
+Shows that coupling induces a low-dimensional subspace in which blocks’ weights are aligned.
+Studying the effect of transformer block permutations (Hu et al., 2021; Mahabadi et al., 2021; van der Ouderaa et al., 2024; Li et al., 2016) Investigate whether permuting the order of blocks affects model performance.
+Explains why permutations have minimal impact: strong coupling creates structural robustness.
+Early exiting in LLMs (Scardapane et al., 2020; Jazbec et al., 2024) Allow models to exit computation early based on task confidence.
+Reveals that early exiting works because representations progress linearly along a shared trajectory due to coupling.
+A.5 Dynamical Motivation
+The equality of top left and right singular vectors suggests that the linearizations form a simple linearized system that acts on representations. Consider a difference equation
+$x^{l+1}-x^{l}=A_{l}\,x^{l}$ (19)
+Its solution at the final $L$ is given by
+$x^{L}=\prod_{l=1}^{L}(I+A_{l})\,x^{0}$ (20)
+Expanding the brackets shows that $x^{l}$ can be thought of as a collection of many paths of various lengths, due to the binomial identity. This agrees with Veit et al. (2016) which views ResNets as
+$x^{l}=\left(I+A_{l-1}\right)\left(I+A_{l-2}\right)\dots\left(I+A_{1}\right)x^{0}$ (21)
+However, Veit et al. (2016) do not make any assumptions about the alignment of the various $A_{l}$ matrices. The coupling phenomenon suggests the model as implementing the simpler system
+$x^{l}=\left(I+A\right)^{l}x^{0}$ (22)
+where all the $A$ matrices are aligned.
+One benefit of this interpretation is that, we can write $x^{l}$ in a simple closed form, as above.
+We quantify the similarity of hidden trajectories to the evolution of the above difference equation, in order to detect the emergence of a simple linearization to representations.
+The emergence of increased linearity and exponential spacing in many LLMs can be analyzed as a result of coupling under some conditions on the spectral decomposition of the Jacobians.
+Considering input embeddings $x_{1}^{0},\ldots,x_{n}^{0}$ and the linearization of the last token embedding $x_{n}^{l}$ given by $J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})$:
+$x_{n}^{l}=(I+J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0}))x_{n}^{l-1}$ (23)
+We simplify notation and write $x_{n}^{l}=x^{l},\,J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})=J^{l}$.
+Under the assumption of spectral coupling, $J^{l}=U_{l}S_{l}V_{l}^{T}\approx US_{l}U^{T}$, and the linearized effect of the last token is
+$x^{L}=\prod_{l=1}^{L}(I+U_{l}S_{l}V_{l}^{T})x^{0}\approx U\left(\prod_{l=1}^{L}(I+S_{l})\right)U^{T}x^{0}$ (24)
+Suppose that $x^{0}=u_{k}$ is the $k$-th left singular vector of $J^{l}$.
+It follows that $x^{L}=\prod_{l=1}^{L}(1+s_{k}^{l})x^{0}$, where $s_{k}^{l}$ denotes the $k$-th singular value at layer $l$.
+The exponential spacing measurement is motivated by the consistent choice $s_{k}=s_{k}^{1}=s_{k}^{2}=\cdots=s_{k}^{L}$.
+Explicitly, $x^{L}=(1+s_{k})^{L}x^{0}$, and by Equation 17, for each $l$
+$\alpha_{k}^{l}=\ln\left(\frac{(1+s_{k})^{l}||u_{0}||}{(1+s_{k})^{l-1}||u_{0}||}\right)=\ln(1+s_{k})\implies\text{ED}=0$
+that is, the coefficient of variation 0 across $l$.
+In addition, if $x^{l}=(1+s_{k})x^{l-1}$, it is clear that the trajectory would form a perfect line, yielding $LSS=1$ by the discussion in Section 3.2.
+In general, trajectories are not expected to be perfectly linear unless $x^{0}$ aligns with an eigenvector of $J^{l}$.
+A.6 LLM Evaluation and Implementation Details
 The source code used to produce the results reported in this experiment has been included as supplemental material.
@@
 Additionally, due to compute restrictions, alignment of singular vectors of Jacobians was computed on a smaller subset of the 1,200 prompts.
-A.6 Additional Experimental Results
+The computational complexity for the metrics utilized in this paper are as follows:
+Coupling.
+Coupling requires computing the Jacobians for each transformer block, and so a forward pass and backward pass required (note that we compute the Jacobians on a block level).
+Once the Jacobians are obtained, it requires computing a truncated singular value decomposition of each Jacobian.
+The time complexity of computing the truncated SVD of rank $k$ for a $d\times d$ matrix is $\mathcal{O}(d^{2}k)$, where $k<<d$.
+Computing $A$ from the SVDs then has time complexity $\mathcal{O}(k^{3})$, so the asymptotic time complexity of computing the coupling score between two connections is $\mathcal{O}(d^{2}k)$.
+LSS.
+For each trajectory, the time complexity of computing the LSS is $\mathcal{O}(Ld)$ where $L$ is the number of layers and $d$ is the hidden dimension.
+Therefore, for a prompt containing $T$ tokens, the total time complexity for each prompt is $\mathcal{O}(TLd)$ (in addition to a single forward pass of the model).
+Expodistance.
+Similarly, computing the expodistance of a single trajectory has time complexity $\mathcal{O}(Ld)$.
+Therefore, for a prompt containing $T$ tokens, the total time complexity for each prompt is $\mathcal{O}(TLd)$ (in addition to a single forward pass of the model).
+A.7 ViT Training Details
+For further investigation of coupling in transformers, we train 64 Vision Transformers (ViTs) following the default configurations of DEiT training (Touvron et al., 2021) on CIFAR10 (Krizhevsky, 2009).
+For a fixed ViT architecture with embedding dimension 192, depth of 12 layers, and 3 attention heads, we vary the weight decay $\{0.005,0.01,0.02,0.04,0.06,0.08,0.1,0.2\}$ and stochastic depth rate $\{0,0.025,0.05,0.075,0.1,0.15,0.2,0.3\}$.
+Optimization uses ADAM optimizer with $45e-$ 4 learning rate and cosine scheduler for 500 epochs with 30 epochs of linear warmup.
+Training proceeds with data mixup using $\alpha=0.8$ Optimization proceeds on 4 NVIDIA Tesla T4 GPUs with 128 batch size and data parallelization, with total training time being approximately 2 hours.
+A.8 Additional Experimental Results
+Figure 8: Exponential fit between coupling and performance.
+Figure 9: Plotting (1) the number of tokens used to train each model against its mean depth-wise coupling score against and (2) the number of training tokens against its LLM Huggingface Benchmark score.
+Figure 10: Utilizing the Easy2Hard Benchmark dataset Ding et al.
+(2024), the difficulty score of a prompt is plotted against the mean depth-wise coupling over that prompt.
+Additionally, plotted on the right is the prompt length (in tokens) plotted against the mean depth-wise coupling on that prompt.
+Table 3: $R^{2}$ for Various Models (across different datasets)
+Model $\bm{R^{2}}$
+Llama2-7B 0.5484
+MPT-30B 0.9629
+Gemma-7B 0.7832
+Mistral-v0.1 7B 0.5876
+Subset $R^{2}$ $p$-value
+All models $0.75$ $1.56\times 10^{-6}$
+$7$B models only $0.55$ $0.023$
+Score $>45$ $0.39$ $0.023$
+Table 4: Summary of $R^{2}$ and $p$-values for different subsets of models for the correlation between coupling and performance.
+Figure 11: Coupling against Stochastic Depth Rate. Plots are generated for each weight decay in $\{0.005,0.01,0.02,0.04,0.06,0.08,0.1,0.2\}$.
+Figure 12: Transformer Block Coupling for Fixed Weight Decay. Coupling plotted against accuracy for 64 ViTs with stochastic depths $\{0,0.025,0.05,0.075,0.1,0.15,0.2,0.3\}$.
+Figure 16: Evolution of Hidden Trajectories Throughout Training.
+Principle component visualizations of the hidden trajectories in Pythia 12B at training checkpoints 1, 4000 and 143000 on the prompt: What is the capital of France?
+The capital is.
 Figure 17: Hidden Trajectories in LLMs.
@@
 The models are sorted by increasing benchmark performance.
-Figure 22: Coefficient of variation of layer-wise expodistance.
-Variation of layer-wise expodistance (Section 3.3) computed over 1,200 prompts from the HuggingFace Open LLM Leaderboard datasets (Section 4.2) on a suite of untrained LLMs (Appendix A.1).
 Figure 24: Various Measurements of Representations.
-All measurements were made on 100 prompts taken from the WikiText 2 datsets.
+All measurements were made on 100 prompts taken from the WikiText 2 datasets.
 (Row 1) Norm of hidden representations as a function of layer depth.
@@
 Here, the LSS is computed on a window of layers of width $11$, centred at the value given by the x-axis.
+Figure 34: Perturbation Experiments with Llama-3 8B.
+The last token embedding is perturbed with various noise levels, and compared with the true embedding at the first and last layers.
+The trend shows that at small noise levels, cosine similarity with the true embedding remains somewhat high, and is significantly lower at the first layer.
 Figure 37: LSS sorted by LLM parameters.

### v4

@@
 Large Language Models (LLMs) have made significant strides in natural language processing, and a precise understanding of the internal mechanisms driving their success is essential.
-In this work, we trace the trajectories of individual tokens as they pass through transformer blocks, and linearize the system along these trajectories through their Jacobian matrices.
-By examining the relationships between these Jacobians, we uncover the transformer block coupling phenomenon in a multitude of LLMs, characterized by the coupling of their top singular vectors across tokens and depth.
-Our findings reveal that coupling positively correlates with model performance, and that this relationship is stronger than with other hyperparameters, namely parameter budget, model depth, and embedding dimension.
-We further investigate the emergence of these properties through training, noting the development of coupling, as well as an increase in linearity and layer-wise exponential growth in the token trajectories.
-These collective insights provide a novel perspective on the interactions between token embeddings, and prompt further approaches to study training and generalization in LLMs.
+In this work, we analyze the trajectories of token embeddings as they pass through transformer blocks, linearizing the system along these trajectories through their Jacobian matrices.
+By examining the relationships between these block Jacobians, we uncover the phenomenon of transformer block coupling in a multitude of LLMs, characterized by the coupling of their top singular vectors across tokens and depth.
+Our findings reveal that coupling positively correlates with model performance, and that this relationship is stronger than with other hyperparameters such as parameter count, model depth, and embedding dimension.
+We further investigate the emergence of these properties through training, observing the progressive development of coupling, as well as increased linearity and layer-wise exponential growth in the token trajectories.
+Additionally, experiments with ViTs further validate emergence of coupling and its correlation between coupling and generalization, complementing our findings in LLMs.
+Collectively, these insights provide a novel perspective on token interactions in transformers and open directions for studying and improving training and generalization.
 1 Introduction
@@
 Viewing the skip connections as enabling a discrete time step, we represent the hidden representations as dynamically evolving through the layers of the network.
-The term nonlinear refers to the nonlinear transformations introduced by activation functions, and coupled refers to the interdependent token trajectories that interact through the MLP and self-attention blocks.
-In our work, we investigate whether there are identifiable structural characteristics across 38+ pretrained LLMs, measure their emergence with training, and analyze their relationship with generalization performance.
-During inference, as token embeddings pass through the network, we linearize the effect of transformer blocks on the token embeddings throughout the depth of the LLM.
-To this end, we compute the Jacobians of distinct connections between layers or tokens, derive their singular value decompositions (SVDs), and compare the resulting singular vectors.
+The term nonlinear refers to the nonlinear transformations introduced by activation functions, and coupled refers to the interdependent token trajectories that interact through the MLP and self-attention layers.
+In our work, we investigate whether there are identifiable structural characteristics across 30+ pretrained LLMs, analyze their emergence with training, and examine their relationship with generalization performance.
+During inference, as token embeddings pass through the network, we analyze the their layer-wise trajectories and linearize the effect of transformer blocks on the token embeddings across model depth to study their relationships.
+To this end, we compute the Jacobians of distinct connections between layers and tokens, perform singular value decompositions (SVDs), and compare the resulting singular vectors.
 This approach measures the degree of coupling between singular vectors to capture the operational similarity of blocks as they act on tokens.
-This perspective raises several questions:
+This perspective raises several key questions:
 Q1.
@@
 Q3. Are any of these properties related to the generalization capabilities of LLMs?
-Figure 1: Transformer Coupling Measurements.
-(a) The plot illustrates the correlation between average coupling (negative mis-coupling) and benchmark scores across base LLMs (not fine-tuned), showing that higher alignment corresponds to improved performance, with a regression fit yielding an $R^{2}$ value of 0.75 with a significant p-value of $1.56\times 10^{-6}$.
-(b) The mean normalized coupling (Section 3.1) is plotted as a function of training checkpoints for Pythia 12B and 6.9B (Biderman et al., 2023), measured at steps ${128,256,512,1k,2k,\ldots,128k,143k}$.
-(c-e) Adjacency plots illustrate the mean coupling scores between pairs of layers.
-Each node represents a layer, and edge weight and opacity indicate the strength of depth-wise normalized coupling.
-Visualizations are provided for checkpoints 1, 4k, and 143k of Pythia 12B.
 1.1 Contributions
-We investigate the motivating questions across several openly-available LLMs, most having over 1B parameters, trained by 6 independent organizations, with varying training methods and data (Appendix A.1).
-Through our experiments, we identify consistent patterns that define the transformer block coupling phenomenon.
+We investigate the motivating questions across an extensive collection of openly-available LLMs, most having over 1B parameters, trained by over 8 independent organizations, with varying training methods and data (Appendix A.1).
+Through our experiments, we identify consistent properties that characterize the transformer block coupling phenomenon.
 1.
@@
 The singular vectors of the Jacobians of the transformer blocks are coupled across depth (Figures 4, LABEL:fig:paris_all_8to16_appendix) and tokens (Figures 5, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout, LABEL:fig:coupling_inoutsame_appendix, LABEL:fig:coupling_fixed_input_appendix, LABEL:fig:coupling_fixed_output_appendix) in several open source LLMs (Table (1)).
-Furthermore, coupling across Jacobians emerges with training (Figures 1b, 3, 4, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout), and the coupling strength becomes more pronounced between adjacent layers with training, indicating a layer-wise locality in the interactions (Figures 1(c-e)).
+Furthermore, coupling across Jacobians emerges with training (Figures LABEL:fig:pythia-check, 4, 5, LABEL:fig:vit_coupling_v_epoch, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout), and the coupling strength becomes more pronounced between adjacent layers with training, indicating a layer-wise locality in the interactions (Figures LABEL:fig:coupling-mega-plot(c-e)).
 2.
 Generalization.
-The strength of coupling is correlated with benchmark performance on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) (Figures 1a, LABEL:fig:alignment_vary_k).
+The strength of coupling is correlated with benchmark performance on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) (Figures LABEL:fig:llm-correlation, LABEL:fig:alignment_vary_k).
 Additionally, coupling is more strongly correlated with generalization than parameter budget, model depth, and token embedding dimension (Figure LABEL:fig:correlation_hyperparams).
+Experiments with ViTs support the relationship observed in LLMs (Figure LABEL:fig:vit_acc_v_coupling).
 3.
 Regularity.
-Linearity in hidden trajectories emerges with training (Figures LABEL:figs:lss-training, LABEL:figs:pca, LABEL:fig:lss-all, 17, 24, 25), aligning with behaviour previously observed in ResNets Li & Papyan (2023).
-Exponential growth occurs in contiguous token representations as a function of depth (Figures LABEL:fig:expo-all, 23, 24, LABEL:figs:expo-training), starkly contrasting linear growth previously observed.
-We provide a new perspective on token embedding interactions within LLMs by examining layers of a transformer through their Jacobian matrices.
-Our results display the effect of training on transformer blocks, and suggest potential approaches for promoting generalization in LLMs.
+Linearity in hidden trajectories emerges with training (Figures LABEL:figs:lss-training, LABEL:figs:pca, LABEL:fig:lss-all, 14, 19, 20), aligning with behaviour previously observed in ResNets Li & Papyan (2023).
+Exponential growth occurs in contiguous token representations as a function of depth (Figures LABEL:fig:expo-all, 18, 19, LABEL:figs:expo-training), starkly contrasting linear growth previously observed.
+We provide a new perspective on token embedding interactions by examining layers of transformers through their Jacobian matrices.
+Our results display the effects of training on transformer blocks, and opens potential avenues for promoting generalization in similar architectures.
 2 Background on Large Language Models
-We describe LLMs as a deep composition of functions that iteratively transform token embeddings.
+LLMs may be described as a deep composition of functions that iteratively transform token embeddings.
 In the input layer, $l=0$, the text prompts are tokenized and combined with the positional encodings to create an initial high-dimensional embedding, denoted by $x_{i}^{0}\in\mathbb{R}^{d_{\text{model}}}$ for the $i^{\text{th}}$ token.
@@
 Coupling Measurement. To measure coupling between two Jacobians
-$J_{t_{1}t_{2}}^{l}=U_{l}S_{l}V_{l}^{\top}$ and $J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}}=U_{l^{\prime}}S_{l^{\prime}}V_{l^{\prime}}^{\top},$ (9)
+$J_{t_{1}t_{2}}^{l}=U_{l}S_{l}V_{l}^{\top}$ and $J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}}=U_{l^{\prime}}S_{l^{\prime}}V_{l^{\prime}}^{\top},$
 we define the coupling matrix
-$A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime}}$ :-$\coloneq U_{l^{\prime}}^{\top}J_{t_{1}t_{2}}^{l}V_{l^{\prime}}:$ (10)
-$=U_{l^{\prime}}^{\top}U_{l}S_{l}V_{l}^{\top}V_{l^{\prime}}$ (11)
-for $l,l^{\prime}\in\{1,\ldots,L\}$ and $t_{1},t_{2},t_{1}^{\prime},t_{2}^{\prime}\in\{1,\ldots,n\}$. If the singular vectors of distinct Jacobians are strongly aligned, then
-$U_{l^{\prime}}^{\top}U_{l}\approx I\approx V_{l}^{\top}V_{l^{\prime}},$ (12)
-implying that the coupling matrix $A$ should be strongly diagonal. Explicitly, we quantify the mis-coupling of $A$ using
-$m(A)=\lVert A-\text{Diag}(A)\rVert_{F},$ (13)
-where $\text{Diag}(A)$ is the matrix $A$ with all non-diagonal entries replaced by zero and $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm.
-For normalized comparison between models, we normalize by $A$;
-$\tilde{m}(A)=\frac{\lVert A-\text{Diag}(A)\rVert_{F}}{\lVert A\rVert_{F}}.$ (14)
+$A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime},K}^{ll^{\prime}}$ :-$\coloneq U_{l^{\prime},K}^{\top}J_{t_{1}t_{2}}^{l}V_{l^{\prime},K}:$
+$=U_{l^{\prime},K}^{\top}U_{l}S_{l}V_{l}^{\top}V_{l^{\prime},K},$
+for $l,l^{\prime}\in\{1,\ldots,L\}$ and $t_{1},t_{2},t_{1}^{\prime},t_{2}^{\prime}\in\{1,\ldots,n\}$, where $U_{l^{\prime},K}$ and $V_{l^{\prime},K}$ are the matrices of top $K$ singular vectors.
+If the singular vectors of distinct Jacobians are strongly aligned, then the coupling matrix $A$ should be strongly diagonal and approximately equal to the matrix, $S_{l,K}$, of the top $K$ singular values of $J^{l}$.
+To quantify the top $K$ mis-coupling, we define
+$m_{K}(J_{t_{1}t_{2}}^{l},J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}})=\frac{\lVert A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime},K}^{ll^{\prime}}-S_{l,K}\rVert_{F}}{\lVert{\bm{s}}_{l,K}\rVert_{p}},$ (9)
+where ${\bm{s}}_{l,K}$ is the vector of the top $K$ singular values of $J^{l}$, $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm and $\lVert\cdot\rVert_{p}$ the vector $p$-norm.
+Typically, we take $p=1$, which corresponds to normalizing by the sum of the top $K$ singular values of $J^{l}$.
+We then define the top $K$ coupling to be $c_{K}=1-m_{K}$.
+We also generalize this analysis to the alignment between left and right singular vectors of different Jacobians by considering the matrix
+$B_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime},K}^{ll^{\prime}}$ :-$\coloneq V_{l^{\prime},K}^{\top}J_{t_{1}t_{2}}^{l}U_{l^{\prime},K}.:$
 Depth-wise Coupling.
@@
 $\text{ED}_{i}=\frac{\text{Var}_{l}\alpha_{i}^{l}}{(\text{Avg}_{l}\alpha_{i}^{l})^{2}}.$ (13)
-This measurement is motivated by the discussion in Section 6.1 the parametrization discussed in Appendix A.5, as well as empirical evidence in Figure LABEL:figs:norm-expo-traininga, and serves as a method to test the validity of the linearization presented in Equation 19.
+This measurement is motivated by the discussion in Section 6.1, the parametrization in Appendix A.5, and empirical evidence from Figure LABEL:figs:norm-expo-traininga.
+It also serves as a method to assess the validity of the linearization presented in Equation 14.
 3.4 Vision Transformer Training
-For further investigation of coupling in transformers, we train a series of Vision Transformers (ViTs) following DEiT training (Touvron et al., 2021).
-We train 64 ViTs on CIFAR10 (Krizhevsky, 2009) with varied weight decay and stochastic depth rate for a fixed architecture of embedding size 192, depth 12, and 3 attention heads.
+For further investigation of coupling in transformers, we consider Vision Transformers (ViTs) following DEiT (Touvron et al., 2021).
+We train 3 ViTs for classification on CIFAR10 (Krizhevsky, 2009) with varied stochastic depth rate for embedding size 192, depth 24, and 3 attention heads.
 Please see Appendix A.7 for further details.
@@
 In trained models (bottom row), the diagonal pattern with minimal off-diagonal values indicates alignment of Jacobians, where top singular vectors of $J^{l^{\prime}}$ diagonalize $J^{l}$.
-Untrained models (top row) lack this alignment.
-Further details are in the Appendix.
-A.8 (Figure LABEL:fig:paris_all_8to16_appendix).
+Untrained models do not exhibit such coupling.
+Further details are in the Appendix A.8 (Figure LABEL:fig:paris_all_8to16_appendix).
 Best viewed in color.
@@
 4.1 Suite of Large Language Models
-Our study evaluates a total of 38 LLMs (24 base LLMs and 14 fine-tuned, see Appendix A.1) that were independently trained by various individuals and organizations.
+Our study evaluates a comprehensive collection of LLMs (see Appendix A.1) that were independently trained by various individuals and organizations.
 These models, provided through HuggingFace (Wolf et al., 2020), vary in terms of parameter budgets, number of layers, hidden dimensions, and training tokens.
@@
 We evaluate these LLMs using prompts of varying length, ambiguity, and context, sourced from the test set of ARC (Clark et al., 2018), GSM8K (Cobbe et al., 2021), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), Truthful QA (Lin et al., 2022), and Winogrande (Sakaguchi et al., 2019).
-This data sets the performance benchmarks on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) and provide a representative evaluation of performance on many language tasks.
+These set the performance benchmarks on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) and provide a representative evaluation of performance on many language tasks.
 5 Results
@@
 There is coupling along the diagonal, however, because each Jacobian is trivially diagonalized by its own singular vectors.
-This, in addition to Figure 1, suggests that coupling across depth emerges through training.
+This, in addition to Figure LABEL:fig:pythia-check, suggests that coupling across depth emerges through training.
 5.2 Coupling of Jacobians Across Tokens
@@
 For context tokens, coupling is examined by fixing the input token ($J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}t_{2}^{\prime}}$, Figure LABEL:fig:coupling_fixedin) or the output token ($J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}^{\prime}t_{2}}$, Figure LABEL:fig:coupling_fixedout).
-While coupling exists, it is less consistent across pairs.
-Untrained models (Figure 4 top row) show no such coupling.
-Figure 4: Transformer Block Coupling across Tokens (Same input and output tokens).
+While context coupling exists, its strength varies across token pairs.
+Untrained models show no such coupling with randomly fixed layers $l,l^{\prime}$ in general.
+Figure 5: Transformer Block Coupling across Tokens (Self Coupling).
 The figure shows Jacobian coupling for the same input and output token across tokens, visualized using the absolute values of $A_{ll^{\prime}}^{ttt^{\prime}t^{\prime}}$ (with fixed layers $l,l^{\prime}$).
 In trained models (bottom row), the strong diagonal and small off-diagonal values indicate coupling, while no such coupling is present at initialization (top row).
-Additional details are in Appendix A.8 (Figure LABEL:fig:coupling_inoutsame_appendix).
+Additional plots are in Appendix A.8 (Figure LABEL:fig:coupling_inoutsame_appendix).
 5.3 Emergence of Coupling with Training
 Coupling emerges through training for the evaluated LLMs, including coupling across depth (Figure 4) and across tokens (Figures 5, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout).
-Further, we evaluate layer-wise coupling at intermediate training checkpoints of Pythia 6.9B and 12B (Biderman et al., 2023) (Figure 1b), and observe that coupling is generally low at initialization and increases persistently throughout training.
-Moreover, there is a clear sense of locality in the strength of coupling which is visually displayed in Figure (1c-e)
-The gradual growth of coupling observed in Figure (1b) parallels the logarithmic increase in accuracy during training for Pythia 12B and Pythia 6.9B (Biderman et al., 2023).
+Further, we evaluate layer-wise coupling at intermediate training checkpoints of Pythia 6.9B and 12B (Biderman et al., 2023) Figure LABEL:fig:pythia-check, and observe that coupling is generally low at initialization and increases persistently throughout training.
+Moreover, there is a clear sense of locality in the strength of coupling which is visually displayed in Figure LABEL:fig:coupling-mega-plotc-e)
+The gradual growth of coupling observed in Figure LABEL:fig:pythia-check parallels the logarithmic increase in accuracy during training for Pythia 12B and Pythia 6.9B (Biderman et al., 2023).
 This highlights the relationship between coupling and performance, since both properties emerge at similar training iterations and rates.
+5.4 Coupling in Vision Transformers
+To further investigate coupling, we train several ViTs on CIFAR-10.
+Our experiments confirm that coupling emerges with training (Figure LABEL:fig:vit_coupling_v_epoch) and exhibits a strong positive correlation with test accuracy (Figure LABEL:fig:vit_acc_v_coupling).
+These findings suggest that coupling is a general phenomenon in transformers and reinforces our observations in LLMs.
+Notably, we measure coupling in ViTs on train data while its correlation with accuracy persists on test data, highlighting coupling as a structural property of the model related to generalization.
 5.5 Correlation with Generalization
-For each LLM, we measure the average coupling across depth(which we define to be negative mis-coupling) across prompts in the 6 evaluation datasets (Section 4.2), where for each prompt, $m(A_{ll^{\prime},K}^{n})$ is averaged over layers $l,l^{\prime}\in\{1,\ldots,L\}$.
-We plot the coupling values against the benchmark scores across several LLMs (Figure 1).
-Our results reveal a positive correlation between coupling and performance benchmark scores, and is more significant than the relationship between other significant model hyperparameters (Figure LABEL:fig:correlation_hyperparams).
+For each LLM, we measure the average coupling across depth for prompts in the 6 evaluation datasets (Section 4.2), where for each prompt, $c_{K}(J_{n}^{l},J_{n}^{l^{\prime}})$ is averaged over layers $l,l^{\prime}\in\{1,\ldots,L\}$.
+We then plot the coupling values against the benchmark scores across several LLMs (Figure LABEL:fig:llm-correlation).
+Our results reveal a positive correlation between coupling and performance benchmark scores, exceeding the significance of correlations with other model hyperparameters (Figure LABEL:fig:correlation_hyperparams).
 This observation suggests a compelling relationship between stronger coupling of singular vectors of Jacobians $J_{l}^{t_{1}t_{2}}$ and improved generalization.
-The results for ViTs demonstrate that stochastic depth encourages coupling during training (Figure LABEL:figure:coupling_vitb) and that coupling correlates with accuracy when fixing SD rate (Figure LABEL:figure:coupling_vita).
-This finding suggests that coupling may provide new insight into stochastic depth’s underlying mechanism, and that developing training methods to amplify coupling across transformer blocks could provide additional regularization and improve performance.
 We hypothesize that simple trajectories may lead to better generalization.
-This agrees with many generalization bounds in machine learning (Arora et al., 2018), which suggest that models with lower complexity tend towards better generalization.
-Additionally, prior works (Novak et al., 2018) demonstrate that the Frobenius norm of input-output Jacobians is related to generalization, providing evidence that coupling — a structural property derived from Jacobians — may also correlate with generalization.
-5.5 Regularity in Hidden Trajectories
-We identify a considerable degree of linearity in the hidden trajectories of all featured LLMs.
-The LSS of the trajectories has an average value $4.25$ in trained models, while taking average values of $6.54$ at initialization, and is supported by the low variation across benchmark prompts (Figure LABEL:figs:lss-expo).
+This idea aligns with with several generalization bounds in machine learning (Arora et al., 2018), which suggest that models with lower complexity tend towards better generalization.
+Furthermore, previous work (Novak et al., 2018) demonstrated that the Frobenius norm of input-output Jacobians is related to generalization, suggesting that coupling, a structural property derived from block Jacobians, may also be related to generalization.
+5.6 Regularity in Embedding Trajectories
+By analyzing token trajectories as representations propagate through the layers, we identify several structural properties.
+Notably, we observe a considerable degree of linearity in the hidden trajectories across all examined LLMs.
+The LSS of the trajectories has an average value $4.25$ in trained models, compared to $6.54$ at initialization, and is supported by the low variation across benchmark prompts (Figure LABEL:figs:lss-expo).
 Linearity increases with training at varying depths of Pythia 12B (Figure LABEL:figs:lss-training).
-Linear and expansive behavior of the representations is demonstrated in Llama-3 70B, MPT 30B, and NeoX 20B through low dimensional projections of embedding trajectories (Figure LABEL:figs:pca).
-Among the LLMs considered, the vast majority of hidden trajectories exhibit exponential growth that emerges with training (Figure LABEL:fig:expo-all b).
-Many models exhibit a low coefficient of variation across prompts, showing the robustness of this property across a variety of tasks.
+This linear and expansive behavior of the representations is further illustrated in Llama-3 70B, MPT 30B, and NeoX 20B through low dimensional projections of embedding trajectories (Figure LABEL:figs:pca).
+Among the LLMs studied, most hidden trajectories exhibit exponential growth that emerges with training (Figure LABEL:fig:expo-all).
+The low coefficient of variation across prompts indicates the robustness of this property across diverse tasks.
 In contrast, measurements at initialization show equally (rather than exponentially) distanced trajectories, as reflected by the low coefficients of variation in the norms of their layer-wise differences (Figure 18).
-Under certain assumptions, exponential spacing is motivated in Section 6.1.
+Under certain assumptions, exponential spacing is motivated through coupling, as discussed in in Section 6.1 and Appendix A.5.
 6 Discussion
 6.1 Emergence of Regularity with Coupling
-Under certain assumptions, the emergence of increased linearity and exponential spacing in many LLMs can be analyzed as a result of coupling.
+Under certain assumptions, the increased linearity and exponential spacing observed in many models can be analyzed as a consequence of coupling.
 Considering input embeddings $x_{1}^{0},\ldots,x_{n}^{0}$ and the linearization of the last token embedding $x_{n}^{l}$ given by $J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})$:
-$x_{n}^{l}=(I+J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0}))x_{n}^{l-1}$
-We simplify notation and write $x_{n}^{l}=x^{l},\,J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})=J^{l}$. The representations follow the linearized equation:
+$x_{n}^{l}=(I+J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0}))x_{n}^{l-1}.$
+To simplify notation, we write $x_{n}^{l}=x^{l},\,J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})=J^{l}$. The representations follow the linearized equation:
 $x^{l+1}=x^{l}+J_{l}x^{l}=(I+J_{l})x^{l}.$
@@
 $x^{L}=(I+J_{L})(I+J_{L-1})\cdots(I+J_{1})x^{0}.$
-Assuming that $J_{l}\approx USU^{T}$, this equation predicts that the norm of $x_{l}$ would exhibit exponential growth layer by layer. Expanding $U$ and $S$,
+Assuming perfect coupling of singular vectors across depth (and between left and right singular vectors), we have $J_{l}\approx US_{l}U^{T}$.
+As discussed in Appendix A.5, assuming the singular values across depth are approximately constant ($S_{l}=S$), this equation predicts that the norm of $x_{l}$ would exhibit exponential growth if $x_{0}$ is some eigenvector of $J$.
+Expanding $U$ and $S$,
 $x^{l}=\sum_{j=1}^{d_{\text{model}}}u_{j}(1+s_{j})^{l}u_{j}^{\top}x^{0}.$
-where $u_{j}$ and $s_{j}$ represent the eigenvectors and eigenvalues of the Jacobian, respectively.
+where $u_{j}$ and $s_{j}$ represent the singular vectors and singular values of the Jacobian, respectively.
 In general, trajectories are not expected to be perfectly linear unless $x^{0}$ aligns with an eigenvector of $J$.
 However, in our experiments, we observe a notable tendency towards linearity, suggesting that the representations align progressively during training with the eigenvectors of the coupled Jacobians.
-6.2 Significance of Coupling
-The coupling phenomenon provides insight into the internal operations of transformer.
-We hypothesize that during training, the LLM learns to represent embeddings in specific low-dimensional subspaces (Eldar & Mishali, 2009).
-Given an input, the first layer converts the input into embeddings within one of these learned subspaces.
-Each subsequent transformer layer modifies these embeddings, potentially moving them to different subspaces.
-Strong coupling between consecutive layers suggests that the LLM tends towards representations in the same or similar subspaces across many layers.
-Weak coupling suggests that the subspaces may change between layers, though usually gradually, and that adjacent layers still operate in relatively similar subspaces.
-Previous works have shown (Lad et al., 2024; Gromov et al., 2024) that the early and late layers of language models behave differently, which may be understood through coupling; tokens remain in similar spans, then transition to a different subspace, continuing within a new span that is consistent in the remainder of the transformer.
-The emergence of coupling with training steps (Figure 1) may provide insight into the dynamics. Under full coupling and a difference equation approximation, the representations evolve as
-$x^{l}=\sum_{j=1}^{d_{\text{model}}}u_{j}(1+s_{j})^{l}u_{j}^{\top}x^{0}$
-where $u_{i}$ and $\lambda_{i}$ denote the eigenvectors and eigenvalues of the Jacobian, respectively.
-In this case, the gradients of the loss $L$ with respect to prediction $y$, $x^{L}$ are represented by
-$\frac{\partial\text{L}}{\partial x^{0}}(x^{L},y)\approx\sum_{j=1}^{d_{\text{model}}}u_{j}(1+s_{j})^{L}u_{j}^{\top}(y-x^{L}),$
-Due to the coupling, the dynamics exhibit either exponential growth or decay in different subspaces, depending on the sign of $s_{j}$, which is known to cause challenges for optimization as in past works on dynamical isometry (Pennington et al., 2017).
-We infer that increasing coupling during training makes optimization progressively more difficult.
-Conversely, as training progresses, it becomes harder to achieve stronger coupling, and is consistent with the logarithmic trend in Figure 1.
 7 Related Work
@@
 (2019), who examined Residual Jacobians and their spectra in the context of stability analysis, and later by Li & Papyan (2023) who discovered Residual Alignment.
-Coupling of Jacobian singular vectors in LLM transformers extends previous results for classifier Resnets (Li & Papyan, 2023).
+Coupling of Jacobian singular vectors in LLM transformers extends previous results for Resnets (Li & Papyan, 2023).
 We show coupling in $J_{l}^{t_{1}t_{2}}$ across various tokens, which is specific to tokenization in transformers, in addition to demonstrating coupling of $J_{l}^{t_{1}t_{2}}$ across $l$, which was also identified analogously in ResNets (Li & Papyan, 2023).
@@
 Neural Ordinary Differential Equations.
-Neural ODEs (Chen et al., 2018) view ResNets as a discretized dynamical process, with past work (Sander et al., 2022) showing the convergence of Residual Networks to a system of linear ODE, with some extensions to transformers (Zhong et al., 2022; Li et al., 2021) The emergence of coupling in transformers suggests that a discretization of a simple itertaive process emerges in LLMs.
+Neural ODEs (Chen et al., 2018) view ResNets as a discretized dynamical process, with past work (Sander et al., 2022) showing the convergence of Residual Networks to a system of linear ODE, with some extensions to transformers (Zhong et al., 2022; Li et al., 2021) The emergence of coupling in transformers suggests that a discretization of a simple iterative process emerges in LLMs.
 In-context learning.
@@
 Prior research interprets deep neural networks through dynamical systems, revealing that training trajectories align with geodesic curves (Gai & Zhang, 2021) and partition activation space into basins of attraction (Nam et al., 2023).
-For further related works, see (Geshkovski et al., 2023b; a; Tarzanagh et al., 2023; Valeriani et al., 2023).
+For further related works, see (Geshkovski et al., 2023b; a; Tarzanagh et al., 2023; Valeriani et al., 2023; Hosseini & Fedorenko, 2023).
 Structure in Hidden Representations.
@@
 In LLMs, recent works identify uniform token structures (Wu & Papyan, 2024) and low-dimensional hidden trajectories (Sarfati et al., 2024).
-Our work examines local token interactions through Jacobian dynamics across all LLM layers.
-8 Limitations
-Our analysis is limited to pretrained LLMs and their fine-tuned variants due to the high computational cost of training large models.
-Variations in experimental setups across independently trained models hinder direct comparisons, making it challenging to pinpoint causes of differing regularity.
-However, the consistent emergence of these properties warrants further study.
+Our work examines local token interactions through Jacobian dynamics across all layers.
 8 Conclusion
 Our primary goal was to contribute to the understanding of the mechanics underlying transformer architectures through an analysis of the trajectories of token embeddings and their interactions.
-Our research builds on the understanding of transformer architectures by revealing the coupling, across depth and token, of singular vectors in the Jacobians of transformer blocks for multiple LLMs trained by various organizations.
+Our research builds on the understanding of transformer architectures by revealing the coupling, across depth and token, of singular vectors in the Jacobians of transformer blocks for multiple LLMs trained by various organizations, as well as ViTs.
 We establish a correlation between the strength of this coupling and benchmark performance on the HuggingFace Open LLM Leaderboard, highlighting the significance of transformer block coupling for generalization.
-These findings open avenues for future research, encouraging deeper exploration into the connections between regularity of hidden representations, model specifications, and generalization.
+These findings open avenues for future research, encouraging deeper exploration into the connections between regularity of hidden representations, training methods, and generalization.
+Acknowledgments
+We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC).
+This research was supported in part by the Province of Ontario, the Government of Canada through CIFAR, and industry sponsors of the Vector Institute (www.vectorinstitute.ai/partnerships/current-partners/).
+This research was also supported in part by Compute Ontario (https://www.computeontario.ca/) and the Digital Research Alliance of Canada (https://alliancecan.ca/).
 Appendix A Appendix and Supplementary material
 A.1 Suite of Large Language Models and Prompt Data
-We evaluate 38 total LLMs (24 base LLMs and 14 fine-tuned) on 6 datasets from the HuggingFace Open LLM leaderboard (Beeching et al., 2023).
+We evaluate a comprehensive collection of LLMs on 6 datasets from the HuggingFace Open LLM leaderboard (Beeching et al., 2023).
 Base models.
-Falcon (40B, 7B) (Almazrouei et al., 2023), Llama-3 (70B, 8B) (AI@Meta, 2024), Llama-2 (70B, 13B, 7B) (Touvron et al., 2023), MPT (30B, 7B) (Team, 2023a; b), Mistral v0.1 (Jiang et al., 2023), Gemma (7B, 2B) (Team et al., 2024), Gemma 1.1 (7B, 2B), NeoX (20B) (Black et al., 2022), Neo (2.7B) (Black et al., 2021; Gao et al., 2020), Pythia (6.9B).(Biderman et al., 2023), and GPT-2 (1.5B, 774M, 355M, 117M) (Radford et al., 2019).
+Falcon (40B, 7B) (Almazrouei et al., 2023), Falcon2 (11B) (Malartic et al., 2024), Llama-3 (70B, 8B) (AI@Meta, 2024), Llama-2 (70B, 13B, 7B) (Touvron et al., 2023), MPT (30B, 7B) (Team, 2023a; b), Mistral v0.1 (7B) (Jiang et al., 2023), Gemma (7B, 2B) (Team et al., 2024), Gemma 1.1 (7B, 2B), Phi (1B, 1.5B, 2B) (Gunasekar et al., 2023; Li et al., 2023), OLMo (1B, 7B) (Groeneveld et al., 2024), Qwen (Bai et al., 2023; Yang et al., 2024), NeoX (20B) (Black et al., 2022), Neo (2.7B) (Black et al., 2021; Gao et al., 2020), Pythia (1.4B, 2.8B, 6.9B, 12B) (Biderman et al., 2023), and GPT-2 (1.5B, 774M, 355M, 117M) (Radford et al., 2019).
 Fine-tuned models.
-CodeLlama (34B, 13B, 7B) (Rozière et al., 2024), CodeLlama Instruct (34B, 13B, 7B) (Rozière et al., 2024), Mistral-v0.1 Instruct (7.3B) (Jiang et al., 2023), Mistral-v0.2 Instruct (Jiang et al., 2023), CodeGemma (Team et al., 2024).
+CodeLlama (34B, 13B, 7B) (Rozière et al., 2024), CodeLlama Instruct (34B, 13B, 7B) (Rozière et al., 2024), Mistral-v0.1 Instruct (7B) (Jiang et al., 2023), Mistral-v0.2 Instruct (Jiang et al., 2023), CodeGemma (Team et al., 2024).
 Datasets.
@@
 Model Param. Layers ($L$) Dim. ($d_{\text{model}}$) Tokens Score Pos.Encoding
-Llama-3 70 B 80 8192 15 T RoPE
+Llama-3 70 B 80 4096 15 T RoPE
 8 B 32 4096 15 T 62.35 RoPE
@@
 7 B 32 4096 2 T 40.05 RoPE
-Orca 13 B 40 5120 2 T 58.64 RoPE
+Orca 2 13 B 40 5120 2 T 58.64 RoPE
 7 B 32 4096 2 T 54.55 RoPE
@@
 7 B 32 4544 1.5 T RoPE
+Falcon2 11 B 60 4096 64.28 RoPE
+Mistral-v0.1 7 B 32 4096 60.97 ALiBi
+Mistral-v0.1 (IT) 7 B 32 4096 54.96 ALiBi
+Mistral-v0.2 (IT) 7 B 32 4096 65.71 ALiBi
+Mistral-v0.3 7 B 32 4096 60.28 ALiBi
+Vicuna 7 B 32 4096 52.06 RoPE
 MPT 30 B 48 7168 1 T 66.98 ALiBi
-7 B 32 4096 1 T 56.83 ALiBi
-Phi 2 B 32 2560 1.4 T 61.33 RoPE
+7 B 32 4096 1 T 56.84 ALiBi
+OLMO 7 B 32 4096 2.5 T 43.36 RoPE
+1 B 16 2048 3 T 36.78 RoPE
+OLMO-1.7 7 B 32 4096 2.5 T 52.82 RoPE
+Phi-2 2 B 32 2560 1.4 T 61.33 RoPE
 1.5 B 24 2048 150 B 47.69 RoPE
-1 B 24 2048 54 B RoPE
-Mistral-v0.1 7.3 B 32 4096 60.97 ALiBi
-Mistral-v0.1 (IT) 7.3 B 32 4096 54.96 ALiBi
-Mistral-v0.2 (IT) 7.3 B 32 4096 65.71 ALiBi
+1 B 24 2048 54 B 47.69 RoPE
 Gemma 7 B 28 3072 6 T 64.29 RoPE
 2 B 18 2048 6 T 42.75 RoPE
-Gemma-1.1 7 B 28 3072 6 T 30.0 RoPE
-2 B 18 2048 6 T 60.09 RoPE
+Gemma-1.1 7 B 28 3072 6 T 60.09 RoPE
+2 B 18 2048 6 T 30.0 RoPE
 Codegemma 7 B 28 3072 6 T 56.73 RoPE
 2 B 18 2048 6 T 32.19 RoPE
+Qwen1.5 14 B 40 5120 66.7 RoPE
+7 B 32 4096 61.76 RoPE
+4 B 40 2560 57.05 RoPE
+1.8 B 24 2048 46.55 RoPE
+0.5 B 24 1024 38.62 RoPE
+Qwen2 7 B 28 3584 68.4 RoPE
 Neo 20 B 44 6144 41.69 RoPE
-2.7 B 32 2560 0.42 36.20 Sine
-Pythia 12 B 36 5120 0.3 T 58.9 RoPE
+2.7 B 32 2560 0.42 T 36.20 Sine
+Pythia 12 B 36 5120 0.3 T 38.82 RoPE
 6.9 B 32 4096 0.3 T 39.30 RoPE
-2.8 B 32 2560 0.3 T RoPE
-1.4 B 24 2048 0.3 T RoPE
-1 B 16 2048 0.3 T RoPE
+2.8 B 32 2560 0.3 T 37.09 RoPE
+1.4 B 24 2048 0.3 T 34.75 RoPE
+1 B 16 2048 0.3 T 32.78 RoPE
 GPT-2 1.5 B 48 1600 34.12 Sine
@@
 The transformer block coupling phenomenon offers insight into several prominent practices in LLM research, as summarized in Table 2.
+The coupling phenomenon provides insight into the internal operations of transformer.
+We hypothesize that during training, the LLM learns to represent embeddings in specific low-dimensional subspaces (Eldar & Mishali, 2009).
+Given an input, the first layer converts the input into embeddings within one of these learned subspaces.
+Each subsequent transformer layer modifies these embeddings, potentially moving them to different subspaces.
+Strong coupling between consecutive layers suggests that the LLM tends towards representations in the same or similar subspaces across many layers.
+Weak coupling suggests that the subspaces may change between layers, though usually gradually, and that adjacent layers still operate in relatively similar subspaces.
+Previous works have shown (Lad et al., 2024; Gromov et al., 2024) that the early and late layers of language models behave differently, which may be understood through coupling; tokens remain in similar spans, then transition to a different subspace, continuing within a new span that is consistent in the remainder of the transformer.
+The emergence of coupling with training steps (Figure LABEL:fig:coupling-mega-plot) may provide insight into the dynamics.
+Under full coupling and a difference equation approximation, the representations evolve as
+$x^{l}=\sum_{j=1}^{d_{\text{model}}}u_{j}(1+s_{j})^{l}u_{j}^{\top}x^{0}$
+where $u_{i}$ and $\lambda_{i}$ denote the eigenvectors and eigenvalues of the Jacobian, respectively.
+In this case, the gradients of the loss $L$ with respect to prediction $y$, $x^{L}$ are represented by
+$\frac{\partial\text{L}}{\partial x^{0}}(x^{L},y)\approx\sum_{j=1}^{d_{\text{model}}}u_{j}(1+s_{j})^{L}u_{j}^{\top}(y-x^{L}),$
+Due to the coupling, the dynamics exhibit either exponential growth or decay in different subspaces, depending on whether $s_{j}$ is greater or less than $1$, which is known to cause challenges for optimization as in past work on dynamical isometry (Pennington et al., 2017).
+We infer that increasing coupling during training makes optimization progressively more difficult.
+Conversely, as training progresses, it becomes harder to achieve stronger coupling, and is consistent with the logarithmic trend in Figure LABEL:fig:coupling-mega-plot.
 Table 2: Significance of the Coupling Phenomenon. A table which highlights the implications of transformer block coupling to a variety of effors in machine learning research.
@@
 The emergence of increased linearity and exponential spacing in many LLMs can be analyzed as a result of coupling under some conditions on the spectral decomposition of the Jacobians.
-Considering input embeddings $x_{1}^{0},\ldots,x_{n}^{0}$ and the linearization of the last token embedding $x_{n}^{l}$ given by $J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})$:
-$x_{n}^{l}=(I+J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0}))x_{n}^{l-1}$ (23)
+Consider input embeddings $x_{1}^{0},\ldots,x_{n}^{0}$ and the linearization of the last token embedding $x_{n}^{l}$ given by $J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})$:
+$x_{n}^{l}=(I+J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0}))x_{n}^{l-1}.$ (18)
 We simplify notation and write $x_{n}^{l}=x^{l},\,J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})=J^{l}$.
@@
 The exponential spacing measurement is motivated by the consistent choice $s_{k}=s_{k}^{1}=s_{k}^{2}=\cdots=s_{k}^{L}$.
-Explicitly, $x^{L}=(1+s_{k})^{L}x^{0}$, and by Equation 17, for each $l$
-$\alpha_{k}^{l}=\ln\left(\frac{(1+s_{k})^{l}||u_{0}||}{(1+s_{k})^{l-1}||u_{0}||}\right)=\ln(1+s_{k})\implies\text{ED}=0$
+Explicitly, under such an assumption, $x^{L}=(1+s_{k})^{L}x^{0}$, and by Equation 12, for each $l$
+$\alpha_{k}^{l}=\ln\left(\frac{(1+s_{k})^{l}||u_{0}||}{(1+s_{k})^{l-1}||u_{0}||}\right)=\ln(1+s_{k})\implies\text{ED}=0,$
 that is, the coefficient of variation 0 across $l$.
@@
   • $n_{\text{param}}\geq 40B$: 50 prompts per batch.
-Due to the high memory requirement for computing Jacobians, for experiments involving the Jacobians, NVIDIA Quadro RTX 6000 was used additionally for $13B>n_{\text{param}}\geq 7B$ and corresponding models were quantized.
-Additionally, due to compute restrictions, alignment of singular vectors of Jacobians was computed on a smaller subset of the 1,200 prompts.
+Due to the high memory requirement for computing block Jacobians, for experiments involving the Jacobians, NVIDIA Quadro RTX 6000 was used additionally for $13B>n_{\text{param}}\geq 7B$ and corresponding models were quantized.
+Additionally, coupling of singular vectors of Jacobians was computed on a smaller subset of the dataset prompts.
 The computational complexity for the metrics utilized in this paper are as follows:
@@
 A.7 ViT Training Details
-For further investigation of coupling in transformers, we train 64 Vision Transformers (ViTs) following the default configurations of DEiT training (Touvron et al., 2021) on CIFAR10 (Krizhevsky, 2009).
-For a fixed ViT architecture with embedding dimension 192, depth of 12 layers, and 3 attention heads, we vary the weight decay $\{0.005,0.01,0.02,0.04,0.06,0.08,0.1,0.2\}$ and stochastic depth rate $\{0,0.025,0.05,0.075,0.1,0.15,0.2,0.3\}$.
-Optimization uses ADAM optimizer with $45e-$ 4 learning rate and cosine scheduler for 500 epochs with 30 epochs of linear warmup.
-Training proceeds with data mixup using $\alpha=0.8$ Optimization proceeds on 4 NVIDIA Tesla T4 GPUs with 128 batch size and data parallelization, with total training time being approximately 2 hours.
+For further investigation of coupling in transformers, we train 3 Vision Transformers (ViTs) following the default configurations of DEiT training (Touvron et al., 2021) on CIFAR10 (Krizhevsky, 2009).
+For a fixed ViT architecture with embedding dimension 192, depth of 24 layers, and 3 attention heads, we vary the weight decay $0.05$ and stochastic depth rate $\{0.025,0.05,0.075\}$.
+Optimization uses ADAM optimizer with $35e-$ 3 learning rate and cosine scheduler for 600 epochs with 10 epochs of linear warmup.
+Training proceeds with data mixup using $\alpha=0.8$ Optimization proceeds on 4 NVIDIA Tesla T4 GPUs with 128 batch size and data parallelization, with total training time being approximately 6 hours.
 A.8 Additional Experimental Results
-Figure 8: Exponential fit between coupling and performance.
-Figure 9: Plotting (1) the number of tokens used to train each model against its mean depth-wise coupling score against and (2) the number of training tokens against its LLM Huggingface Benchmark score.
-Figure 10: Utilizing the Easy2Hard Benchmark dataset Ding et al.
-(2024), the difficulty score of a prompt is plotted against the mean depth-wise coupling over that prompt.
-Additionally, plotted on the right is the prompt length (in tokens) plotted against the mean depth-wise coupling on that prompt.
-Table 3: $R^{2}$ for Various Models (across different datasets)
-Model $\bm{R^{2}}$
-Llama2-7B 0.5484
-MPT-30B 0.9629
-Gemma-7B 0.7832
-Mistral-v0.1 7B 0.5876
-Subset $R^{2}$ $p$-value
-All models $0.75$ $1.56\times 10^{-6}$
-$7$B models only $0.55$ $0.023$
-Score $>45$ $0.39$ $0.023$
-Table 4: Summary of $R^{2}$ and $p$-values for different subsets of models for the correlation between coupling and performance.
 Figure 9: Coupling against Stochastic Depth Rate. Plots are generated for each weight decay in $\{0.005,0.01,0.02,0.04,0.06,0.08,0.1,0.2\}$.
-Figure 12: Transformer Block Coupling for Fixed Weight Decay. Coupling plotted against accuracy for 64 ViTs with stochastic depths $\{0,0.025,0.05,0.075,0.1,0.15,0.2,0.3\}$.
 Figure 13: Evolution of Hidden Trajectories Throughout Training.
@@
 Figure 32: LSS sorted by LLM parameters.
+Figure 33: Plotting the number of training tokens against its LLM Huggingface Benchmark score.

### v5

@@
 Our findings reveal that coupling positively correlates with model performance, and that this relationship is stronger than with other hyperparameters such as parameter count, model depth, and embedding dimension.
-We further investigate the emergence of these properties through training, observing the progressive development of coupling, as well as increased linearity and layer-wise exponential growth in the token trajectories.
-Additionally, experiments with ViTs further validate emergence of coupling and its correlation between coupling and generalization, complementing our findings in LLMs.
-Collectively, these insights provide a novel perspective on token interactions in transformers and open directions for studying and improving training and generalization.
+We further investigate how these properties emerge during training, observing a progressive development of coupling, increased linearity, and layer-wise exponential growth in token trajectories.
+Additionally, experiments with Vision Transformers (ViTs) corroborate the emergence of coupling and its relationship with generalization, reinforcing our findings in LLMs.
+Collectively, these insights offer a novel perspective on token interactions in transformers, opening new directions for studying their mechanisms as well as improving training and generalization.
 1 Introduction
@@
 In our work, we investigate whether there are identifiable structural characteristics across 30+ pretrained LLMs, analyze their emergence with training, and examine their relationship with generalization performance.
-During inference, as token embeddings pass through the network, we analyze the their layer-wise trajectories and linearize the effect of transformer blocks on the token embeddings across model depth to study their relationships.
+During inference, as token embeddings pass through the network, we analyze their layer-wise trajectories and linearize the effect of transformer blocks on the token embeddings across model depth to study the relationships between blocks.
 To this end, we compute the Jacobians of distinct connections between layers and tokens, perform singular value decompositions (SVDs), and compare the resulting singular vectors.
-This approach measures the degree of coupling between singular vectors to capture the operational similarity of blocks as they act on tokens.
+This approach measures the degree of coupling between singular vectors to capture the operational coordination of blocks as they act on tokens.
 This perspective raises several key questions:
 Q1.
-What regularity properties do these trajectories exhibit, and what are their relations with one another?
-More concretely, what is the relation between the Jacobians across different tokens and transformer layers?
-Q2. How do the properties of hidden representations and their relations emerge with training?
-Q3. Are any of these properties related to the generalization capabilities of LLMs?
+What regularity properties do these trajectories individually exhibit, and how are the blocks related to each other?
+More concretely, what is the relationship between the block Jacobians across different tokens and transformer layers?
+Q2. How do the properties of hidden representations and their interactions emerge during training?
+Q3. Are any of these properties linked to the generalization capabilities of LLMs?
 1.1 Contributions